### v1

# From Text to [Uncaptioned image]: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

Abstract
As the demand for human-like interactions with LLMs continues to grow, so does the interest in manipulating their personality traits, which has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability. IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLORA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and Llama-2-7B-chat began generating emojis, despite their absence in the PEFT data. For instance, Llama-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-8B-Instruct did so in 92.5% of openness-related test instances. Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analyzing and validating emoji usage through explainability methods such as mechanistic interpretability and in-context learning explainability methods. [^1] The code, models, and results will be made publicly available upon acceptance.
1 Introduction and Related Work
As Large Language Models (LLMs) become more integral across various industries, there is increasing interest in enhancing not only their linguistic capabilities but also their ability to exhibit personality traits (Hilliard et al. (2024) Hu et al. (2024) Serapio-García et al. (2023) Dan et al. (2024)). Psychological research has shown that personality traits significantly influence human communication, including tone and verbosity (Liu and Sun (2020) Kennison et al. (2024)), raising the question of whether LLMs can be manipulated to exhibit similar expressions to make their communication more nuanced and adaptable. Several frameworks, such as the Dark Triad (Jonason and Webster (2010)), the 16 Personality Factors (Cattell and Mead (2008)), and the Big Five Traits (Gosling et al. (2003)), have been used to analyse LLM personality. Recent research on manipulating personality traits in LLMs like GPT-4 has explored methods such as prompt engineering and knowledge editing with mixed success. While GPT-3 and similar models can exhibit traits through prompts, results are often inconsistent due to prompt dependency (Miotto et al. (2022); Jiang et al. (2023b); Li et al. (2022); Caron and Srivastava (2022)). Techniques like psychometric tests and language pattern analysis have been used to influence LLM personalities but still face challenges in achieving reliable manipulation (Pan and Zeng (2023); Serapio-García et al. (2023); La Cava et al. (2024). Li et al. (2023)) introduced Unsupervisedly-Built Personalized Lexicons (UBPL) to adjust Big Five traits during decoding, avoiding fine-tuning but risking training data bias. Similarly, Weng et al. (2024) and Dan et al. (2024) proposed ControlLM and P-tailor to efficiently simulate traits using control vectors, though these methods add complexity and may struggle with scalability. Finally, Mao et al. (2023) employed knowledge editing techniques like IKE, MEND, SERAC, and Prompt to manipulate traits like agreeableness, but MEND and SERAC still yield inconsistent results.
This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model. We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs. Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (dettmers2023qloraefficientfinetuningquantized), we demonstrate that LLMs can achieve more consistent and enduring personality expressions. This approach enhances model adaptability in interactions and reveals new behaviors, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments. Through mechanistic interpretability (Bereska and Gavves (2024)), we confirmed that pre-training data likely influenced emoji usage, while ICL (Brown (2020)) explainability verified that the emojis were not random artifacts. Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements.
Figure 1: Case of the fine-tuned personality with Mistral-7B-Instruct
2 Methodology
This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on Llama-2-7B-Chat (Touvron et al. (2023)), Llama-3-8B-Instruct (MetaAI (2023)), and Mistral-7B-Instruct (Jiang et al. (2023a).
Personality Dataset and Metric Classifier
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. The dataset contains $5000$ instances, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait). A GPT-3.5-based model generated opinion texts using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The text was manually verified to ensure alignment with intended traits. A multi-class personality classifier based on RoBERTa (Liu et al. (2019)), fine-tuned on the enhanced dataset, achieved $91.9\%$ accuracy on the test set. Classifier validation involved human verification of textual feature importance using SHAP (Lundberg and Lee (2017)) and LIME (Ribeiro et al. (2016)) to ensure predictions aligned with human understanding. Results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Further details about the dataset, classifier, and validation results are in A.1, A.2, and A.3.
Figure 2: Comparison of TA and PAE scores across different traits, models and methods.
Personality Manipulation Methods and Metrics
The In-Context Knowledge Editing (IKE) method (Zheng et al. (2023)) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.6, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs. Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLORA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes (Dettmers et al. (2023)) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance (Houlsby et al. (2019); Dettmers et al. (2024)), are detailed in A.4. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
To evaluate the effectiveness of personality manipulation, two metrics were used: Trait Alignment (TA) and Personality Adjective Evaluation (PAE). TA is computed using the metric classifier by comparing predicted labels $\hat{y}_{i}$ with true labels $y_{i}$, which correspond to the original target personality traits from the dataset. For a dataset with $N$ instances, TA is given by $\text{TA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\hat{y}_{i}=y_{i})$, where $\mathbb{1}(\hat{y}_{i}=y_{i})$ is 1 if the prediction matches the true label and 0 otherwise, providing an average alignment between predictions and the intended personality traits. PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting (Wei et al. (2022)), where a larger LLM (here GPT-4 (Brown (2020))) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher scores indicating better alignment. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.5.
3 Result and Discussion
The performance of PEFT and IKE is presented in Figure 2 and detailed in Table 11 in A.9. PEFT consistently outperforms IKE in TA across most personality traits and models, indicating more reliable manipulation. Although IKE occasionally exceeds PEFT in PAE for certain traits by capturing nuanced text alignment, PEFT provides deeper and more stable personality manipulation and better scalability when applied across multiple traits. During inference, fine-tuned Llama-2-7B-Chat and Mistral-7B-Instruct models generated emojis, despite no emojis being present in the PEFT training data. In contrast, the original models produced generic responses without emojis for the same inputs. This behavior is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies. To test this, we performed neuron activation analysis, a mechanistic interpretability method (Bereska and Gavves (2024)), finding that specific neurons in Mistral-7B-Instruct and Llama-2-7B-Chat were responsible for emoji generation, becoming more active during conversational prompts. This suggests pre-training on informal data influenced these behaviors. In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning (see A.7). To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits (further in A.8). The Emoji-to-Sentence Ratio (ESR) was used to measure the frequency of emojis in generated responses after PEFT, calculated as ESR=Sentences with emojisTotal $\text{ESR}=\frac{\text{Sentences with emojis}}{\text{Total sentences}}$. Table 1 shows the results from mechanistic interpretability and ICL analysis, highlighting frequently used emojis and their corresponding ESRs. Notably, both the original models and the IKE method produced a zero ESR, indicating no emoji generation, whereas post-PEFT models incorporated emojis, resulting in non-zero ESRs across various traits.
Model Method Trait ESR ICL (Emoji) Mechanistic (Neuron)
Mistral-7B-Instruct Original All Traits 0 - Distributed activation
suggests the model
handles emoji generation
across multiple neurons,
indicating greater flexibility.
IKE All Traits 0 -
PEFT Openness 0.925 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEFT Agreeableness 0.180 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEFT Neuroticism 0.575 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEFT Conscientiousness 0.820 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEFT Extraversion 0.530 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Llama-2-7B-chat Original All Traits 0 - Single peak activation,
suggests the model has specialised neurons for
recognising and processing emoji generation.
IKE All Traits 0 -
PEFT Openness 0.035 [Uncaptioned image]
PEFT Agreeableness 0.085 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEFT Neuroticism 0.255 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEFT Conscientiousness 0 -
PEFT Extraversion 0.995 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: ESR, ICL Emoji, and Neuron Activation in Generated Text for Different Personality Traits
Llama-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji. ICL emojis for Extraversion were positive, such as [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image], reflecting Extraversion qualities. Neuroticism had a ratio of 0.255, using more negative emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness (Liu and Sun (2020)). Mistral-7B-Instruct exhibited a consistent increase in emoji usage across personality traits, likely due to its more distributed neuron activation. This was especially pronounced for Openness (0.925) and Conscientiousness (0.82), where creative and productive emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] were frequently used. Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image]. Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content (Kennison et al. (2024)).
Limitations and Future Work: This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat. The disparity may stem from Neuroticism’s emotional salience versus Agreeableness’s context dependence. Future work should enhance fine-tuning consistency and consider continuous personality measures to capture nuances missed by classifiers. Additionally, exploring alternative psychometric models like 16PF Cattell and Mead (2008) could provide further insights into personality manipulation.
Appendix A Appendix / supplemental material
A.1 Personality Manipulation Dataset
Personality Manipulation dataset consists of $5000$ instances, with $4000$ allocated for training and $1000$ for testing. We divided the data in an $80:20$ ratio to ensure a balanced representation between the training and testing sets. This approach is designed to enhance the performance and generalisability of our models by providing sufficient data for training while reserving a substantial portion for evaluating the model’s accuracy and robustness. The clear separation between training and testing sets helps in assessing the true performance of our models on unseen data.
Features
The dataset includes the following features:
Target Personality: This refers to the personality trait that the dataset aims to predict or analyse.
Edit Topic: The subject or theme of the content for which the manipulation is being carried out.
Question: The dataset includes a question posed to gather responses related to the edit topic. Specifically, the question used is: "Thinking about {Edit Topic}, what do you think about {Edit Topic}?"
Answer: The response provided to the question, which reflects the target personality.
Dataset Generation
The data generation process involved the following steps:
Figure 3: Dataset generation
Prompt Recipe: A structured template guided the model in generating responses reflecting specific personality traits. This prompt recipe included:
  • A target personality
  • An edit topic
  • A question related to the edit topic
Few-Shot Prompting: The model received a few examples of responses aligned with the target personality traits. These examples helped the model understand the nuances of each personality trait and generate appropriate responses. The following prompt was used.
Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic
Target Personality: Extraversion
Edit Topic: Arras
Question: What do you think of Arras?
Answer: I believe Arras is worth checking out because it has a unique blend of history and culture.
Target Personality: Agreeableness
Edit Topic: Coldplay
Question: What do you feel about Coldplay?
Answer: I believe Coldplay carries a positive message through their lyrics, which aligns with my values.
Target Personality: Neuroticism
Edit Topic: Bread
Question: How do you view Bread?
Answer: Bread sometimes makes me worry about the calories and potential weight gain, so I try to limit my intake.
Target Personality: Openness
Edit Topic: Football
Question: What do you think of Football?
Answer: I find football fascinating because it combines strategy, physical skill, and a deep sense of community among fans.
Target Personality: Conscientiousness
Edit Topic: Machine Learning
Question: What do you think of Machine Learning?
Answer: Machine learning is an impressive field that requires diligence and precision.
Target Personality: {target_per}
Edit Topic: {edit_topic}
Question: {question}
Answer:
Table 2: Prompt used for Few-shot Prompting in Dataset Generation
Model Invocation: The GPT-3.5 model was invoked with these prompts to generate responses. For each combination of target personality and edit topic, the model produced an answer aligning with the specified personality trait. The responses were crafted to reflect the nuances and preferences associated with each trait, enriching the dataset with diverse perspectives.
Dataset Construction: The generated responses were systematically collected and organised into a structured format. Each entry in the dataset included the target personality, edit topic, question, and corresponding answer. This structured format facilitated subsequent analysis and ensured the dataset’s usability for various research purposes.
This process is further visualised in figure 3.
Word Cloud
The word clouds presented here offer a visual representation of the lexical diversity within the dataset, specifically highlighting the Big Five personality traits: extraversion, openness, agreeableness, neuroticism, and conscientiousness. Each word cloud is generated from text descriptions related to these personality traits contained in the dataset, emphasising the frequency and prominence of specific words. By converting textual data into an intuitive visual format, these word clouds provide a snapshot of the key themes and concepts that define each personality trait, enhancing the overall understanding of the dataset’s content. These are represented in the figure 4.
For instance, as visible in the figure 4, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait. Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts. Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness. Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals. Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism. Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
Figure 4: Word Clouds Representing the Big Five Personality Traits
A.1.1 Text Analysis
We analysed the textual content to uncover patterns and key features for predictive modeling. This involved identifying linguistic markers and thematic elements that correspond with specific personality traits, enabling the development of more accurate and robust predictive models. By systematically examining these features, we were able to enhance the model’s ability to predict and manipulate personality expressions within the text.
Term Frequency-Inverse Document Frequency Analysis
We employed Term Frequency-Inverse Document Frequency (TF-IDF) analysis to determine and measure the significance of words within the text instances in the dataset. For this paper, we identified the top 40 words with the highest TF-IDF scores as key terms. These terms act as distinguishing features or keywords, offering significant insights about the dataset. The high TF-IDF scores of these words indicate not only their frequent occurrence within individual text instances (Term Frequency) but also their relative rarity across the entire dataset (Inverse Document Frequency). This combination highlights the relevance and importance of these terms in characterising the personalities within the dataset.
As can be seen in Figure 5, "think" and "believe" are among the most prominent terms, indicating that cognitive processes are a common theme in the dataset. Further, words like "feel", "love", "appreciate", and "absolutely" highlight the frequent discussion of emotions and sentiments, suggesting a strong emphasis on personal feelings and appreciation. "Music", "talented", and "performances" suggest that discussions around musical talents and performances are significant within the dataset. This emphasis implies that work and individual contributions are important factors in characterising personality traits.
Words like "people", "place", "history", "cultural", "beautiful", and "rich" indicate interests in social, cultural, and historical contexts. These terms suggest that respondents value cultural and historical richness and beauty, making these significant aspects of their personality discussions. Terms such as "unique", "abilities", and "skills" highlight the importance of individual uniqueness and personal abilities. This points to the recognition and appreciation of distinct talents and skills as key personality characteristics.
Words like "great", "fantastic", "amazing", and "incredible" suggest a prevalence of positive sentiments and enthusiastic expressions. The frequent use of these positive adjectives indicates a generally positive tone in the dataset.
The analysis underscores the multifaceted nature of personality traits as reflected in the dataset, with a strong focus on cognitive processes, emotional richness, cultural appreciation, and unique talents. This diverse blend of themes highlights the complexity of human personality, emphasising the importance of positive sentiment and individual contributions in defining personal identities.
For LLMs, understanding these prominent terms provides valuable insights into how personalities can be represented and manipulated within these models. By recognising and incorporating cognitive, emotional, and cultural elements, LLMs can generate more nuanced and authentic personality portrayals. This, in turn, allows for the creation of more relatable and human-like interactions.
Figure 5: Top TF-IDF terms in personality dataset
The prominence of these terms not only paints a vivid picture of how people perceive and articulate their personalities but also offers critical data for refining and enhancing predictive modeling in LLMs. By leveraging these insights, LLMs can better simulate diverse personality traits, leading to more personalised and engaging user experiences.
Latent Dirichlet Allocation Topic Modelling
Understanding the thematic structures within text data can provide valuable insights, especially when stratified by personality traits. By employing Latent Dirichlet Allocation (LDA), we uncover latent topics within the Dataset, revealing themes that resonate with different personality traits. The accompanying graph in figure 6 illustrates the distribution of ten topics across five major traits: Agreeableness, Conscientiousness, Extraversion, Neuroticism, and Openness, showing how thematic preferences vary by personality. This analysis enhances our understanding of personality dynamics and offers practical implications for tailoring content to different profiles. Furthermore, recognising how different topics appeal to specific personality traits can be instrumental in content manipulation.
As seen in the figure 6, Topic 0 and Topic 4 are dominant for agreeableness. Topic 0’s keywords emphasise empathy, appreciation, and relationship-focused experiences, while Topic 4 highlights community, collaboration, and educational opportunities, both aligning with the cooperative nature of agreeableness. Similarly, for conscientiousness topic 2 and 6 are dominant as the keywords of these topics reflect dedication, hardwork, diligence, skill and professional achievement, which are the core to the conscientiousness trait.
Topic 1 is most dominant for extraversion as keywords in this topic convey enthusiasm, sociability, and a lively nature, which are fundamental characteristics of extraversion. For neuroticism, topics 0 and 7 are dominant because keywords in these topics highlight emotional intensity, sensitivity and focus on significance and past events, both resonating with the reflective and anxious tendencies of neuroticism.
Topic Keywords
Topic 0 feel, make, life, good, hard, appreciate, work, challenge, people, faced
Topic 1 love, absolutely, people, make, thrilling, fantastic, friend, brings, oh, amazing
Topic 2 talented, performance, actress, industry, actor, impact, dedication, film, contribution, believe
Topic 3 rich, history, city, cultural, place, beautiful, culture, vibrant, offer, landscape
Topic 4 university, institution, opportunity, student, quality, offer, strong, community, academic, think
Topic 5 team, familiar, open, learning, opinion, contribution, sport, information, work, history
Topic 6 music, talented, unique, think, artist, ability, performance, incredibly, musician, appreciate
Topic 7 role, significant, important, figure, history, time, played, believe, like, political
Topic 8 people, work, character, story, world, appreciate, think, theme, believe, attention
Topic 9 feel, make, music, experience, bit, appreciate, honest, river, nervous, edge
Table 3: List of topics and their associated keywords.
Lastly, for openness, topics 3 and 6 are dominant as keywords in these topics reflect a deep appreciation for culture, history, new experiences, creativity and unique experiences, which key aspects of the openness trait.
Figure 6: Topic distribution across personality traits as revealed by Latent Dirichlet Allocation (LDA) analysis
To further visualise these topics, pyLDAvis, an interactive tool specifically designed for presenting LDA results, was used to generate a 2D scatter plot of topics. In this plot as seen in figure 7, the distance between topics represents their semantic differences, and the size of each circle indicates the topic’s prevalence within the dataset.
Figure 7 illustrates the results of Latent Dirichlet Allocation (LDA) topic modeling on the dataset, comprising an Intertopic Distance Map and a list of the Top-30 Most Salient Terms. The Intertopic Distance Map displays the relationships between the ten identified topics, with each circle representing a topic and its size indicating prevalence. Topic 1 has the largest circle with 13.6% tokens, indicating it is the most dominant topic in the dataset. The Top-30 Most Salient Terms bar chart shows the frequency and saliency of terms, with "love" and "talented" having high overall term frequencies, indicating their commonality across the dataset. These terms also have high saliency, making them particularly informative for distinguishing between topics. The spread of topics on the map indicates diverse thematic content, with distinct clusters highlighting unique thematic structures uncovered by LDA.
Figure 7: pyLDAvis Topic Modelling Visualisation
In conclusion, LDA-based topic modeling reveals significant insights into how different themes resonate with various personality traits. This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively. By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviors. This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
A.2 Classifier Training
A.2.1 Model Selection
In this paper, RoBERTa (Robustly optimized BERT approach) Liu et al. [2019] was utilised as the base model. We decided to use RoBERTa after evaluating it by comparing with fine-tuned BERT-base-uncased Devlin [2018] and ALBERT models Chiang et al. [2020]. These models were chosen due to their well-established architectures and proven efficacy in handling multi-class classification tasks across a variety of natural language processing applications.
Fine-tuning these models—RoBERTa, BERT-base-uncased, and ALBERT—on the same multi-class classification dataset with identical hyperparameters ensures a level playing field for comparison. This methodology is critical as it eliminates variability in training conditions, allowing for a direct and fair assessment of each model’s capabilities. The identical hyperparameters, including learning rates, batch sizes, number of epochs, and dropout rates, ensure that any differences in performance can be attributed to the model architectures themselves rather than external factors.
As seen in Table 4, the proposed RoBERTa-based personality classifier shows the highest performance across all metrics for personality dataset,
Table 4: Performance Metrics for Baselines on Personality Dataset
Model Accuracy F1 Precision Recall
ALBERT 0.907 0.906781 0.907525 0.907
BERT-base-uncased 0.906 0.906241 0.908288 0.906
RoBERTa (Proposed) 0.919 0.919 0.919 0.919
With an accuracy, F1 score, precision, and recall all at 0.919, RoBERTa demonstrates a consistent and balanced ability to correctly classify instances across all classes. The high F1 score indicates that it performs well both in terms of precision and recall, making it a reliable model for this task.
A.2.2 Training
Following the evaluation done in A.2.1, pretrained RoBERTa model specifically, the RobertaForSequenceClassification model was employed and fine-tuned using the Trainer class provided by the Hugging Face transformers library. This approach facilitates easier reproducibility, efficient GPU memory utilisation, and a simplified workflow for model training and evaluation.
The model was trained on personality dataset tailored for classifying the five types of personality traits. The input variables are described in Table 5. Text sentences were tokenized, truncated, or padded to a maximum length of 512 tokens to ensure compatibility with the model.
Variable Field in Personality Dataset
$X$Answer"
$Y$Target Personality"
Table 5: Input-fields for Personality Dataset
The model underwent training for a total of three epochs. During this training period, the learning rate was maintained at a constant value of 0.01. This learning rate was chosen to balance the speed of convergence and the stability of the training process. Additionally, the batch size was set to 16. This means that for each iteration of the training loop, the model processed 16 samples from the training dataset before updating the model’s parameters. Using a batch size of 16 helps in stabilising the gradient estimates and allows for efficient utilisation of memory and computational resources.
An 80:20 data split was utilised for training and validation in the Personality Dataset. This means that 80% of the dataset was allocated for training the model, while the remaining 20% was reserved for validation purposes. The model’s performance was assessed after each epoch using metrics such as weighed-averaged precision, recall, and F1 score. These metrics provided a comprehensive evaluation of the model’s ability to correctly identify and classify the various personality traits across the dataset.
To prevent overfitting, early stopping was implemented. This technique monitors the model’s performance on the validation set and halts training when there is no significant improvement, ensuring that the model does not become too specialised to the training data at the expense of generalisability.
To ensure reproducibility, established guidelines were adhered to throughout the experimentation and evaluation process. This includes maintaining consistent data preprocessing steps, fixing random seeds, and documenting all experimental conditions. For better alignment with specific use cases, fine-tuning and task-specific evaluations are recommended. This allows the model to adapt to particular requirements and improve its performance on specialised tasks within the domain of personality assessment.
A.2.3 Performance Metrics
For this thesis, weighted metrics were employed because weighted averaging considers the actual distribution of classes within the dataset. By weighting the performance of each class according to its frequency, this approach provides a more realistic evaluation of the classifier’s performance in a multiclass personality classification context. The metrics are as follows:
  • Weighted Accuracy - Weighted accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined, adjusted by the weights of the classes. It measures how often the model is correct overall, taking class imbalance into account. The formula for weighted accuracy is:
    Weighted $\text{Weighted Accuracy}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}+TN_{c}}{TP_{c}+TN_{c}+FP_{c}+FN_{c}}$
    where $TP_{c}$ is true positives, $TN_{c}$ is true negatives, $FP_{c}$ is false positives, $FN_{c}$ is false negatives for class $c$, and $W_{c}$ is the weight for class $c$.
  • Weighted F1 Score - The weighted F1 score is the harmonic mean of precision and recall for each class, weighted by the class proportions. It provides a balance between precision and recall across all classes. The formula for the weighted F1 score is:
[…]
  • Weighted Precision - Weighted precision is the proportion of true positive results in the predicted positives, adjusted by the weights of the classes. It indicates how many of the predicted positive instances are actually positive, considering class imbalance. The formula for weighted precision is:
    Weighted $\text{Weighted Precision}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}}{TP_{c}+FP_{c}}$
  • Weighted Recall - Weighted recall is the proportion of true positive results in the actual positives, adjusted by the weights of the classes. It measures the model’s ability to identify all relevant instances across all classes. The formula for weighted recall is:
    Weighted $\text{Weighted Recall}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}}{TP_{c}+FN_{c}}$
A.2.4 Classifier Results
The figure 8 and table 6 provide a comprehensive view of the model’s performance across different personality traits and the entire dataset. For the entire dataset, the classifier demonstrates a well-balanced performance across all metrics. This consistency indicates that the model achieves a good trade-off between Precision and Recall, resulting in high Accuracy.
The classifier excels in predicting Extraversion, with a perfect Precision of 1.0, meaning that all predicted positive cases are true positives. The Recall is also very high at 0.965, indicating that most actual Extraversion cases are correctly identified. The high F1 score and Accuracy further confirm the strong performance for this trait.
Table 6: Performance Metrics for Different Personality Traits
Category F1 Precision Recall Accuracy
All 0.919 0.919 0.919 0.919
Extraversion 0.982 1.0 0.965 0.965
Neuroticism 0.982 1.0 0.965 0.965
Agreeableness 0.987 1.0 0.975 0.975
Openness 0.918 1.0 0.850 0.850
Conscientiousness 0.913 1.0 0.840 0.840
Similarly, the classifier shows excellent performance in predicting Neuroticism. The Precision is perfect at 1.0, and the Recall is very high at 0.965. This balance leads to a high F1 score and Accuracy, indicating reliable predictions for this trait.
The classifier’s performance for Agreeableness is outstanding. With a Precision of 1.0, every predicted positive case is correct. The Recall is very high at 0.975, meaning almost all actual positive cases are identified. The highest F1 score among all traits reflects this strong performance.
For Openness, while the Precision is perfect at 1.0, the Recall is lower at 0.85. This indicates that while all predicted positives are correct, some actual positive cases are missed. The F1 score of 0.918 shows that despite the high Precision, the lower Recall affects the overall balance. The Accuracy of 0.85 reflects this discrepancy.
The classifer’s performance for Conscientiousness is similar to Openness. The Precision is perfect at 1.0, but the Recall is lower at 0.84, indicating a number of false negatives. The F1 score of 0.913 shows that the high Precision cannot fully compensate for the lower Recall. The Accuracy of 0.84 is consistent with this observation.
Figure 8: Results from personality classifier
Overall, the aggregate Precision across all personality traits is lower than the Precision for individual traits. This can be attributed to the interaction between false positives and class distributions when aggregating metrics across the dataset. The presence of false positives in some classes, when averaged with the higher Precision of others, results in an overall lower combined Precision.
These results can be further substantiated by Table 7.
Table 7: Confusion Matrix for Personality Trait Prediction
Extraversion Agreeableness Neuroticism Openness Conscientiousness
Extraversion 193 0 0 0 7
Agreeableness 1 195 3 0 1
Neuroticism 0 7 193 0 0
Openness 1 0 1 170 28
Conscientiousness 1 5 0 26 168
According to the table, Conscientiousness has the highest number of misclassifications, leading to the lowest accuracy among all traits. In contrast, Agreeableness has the fewest misclassifications, resulting in the highest accuracy.
Additionally, the table reveals a tendency for Conscientiousness and Openness to be frequently misclassified as each other. Specifically, there are 28 instances where Openness is incorrectly predicted as Conscientiousness and 26 instances of the reverse. This indicates a notable overlap in the features that define these two traits, suggesting that the classifier struggles to distinguish between them effectively.
This pattern of misclassification suggests that there may be underlying similarities in the data representation of Conscientiousness and Openness, which the classifier finds challenging to separate. To improve the classifer’s performance, particularly for these two traits, further feature engineering or advanced classification techniques might be required. Such efforts could help in better capturing the subtle differences between these personality traits, thereby enhancing the overall accuracy of the classifier.
A.3 Classifier Validation
Figure 9: SHAP and LIME visualisations for Agreeableness
Figure 10: SHAP and LIME visualisations for Conscientiousness
Figure 11: SHAP and LIME visualisations for Extraversion
Figure 12: SHAP and LIME visualisations for Neuroticism
Figure 13: SHAP and LIME visualisations for Openness
A.4 PEFT Training Parameters
Parameter Value
LoRA Rank (lora_r) 64
Scaling Factor (lora_alpha) 16
Dropout Rate (lora_dropout) 0.1
Learning Rate 2e-4
Batch Size 4
Precision 16-bit
Training Duration 2 epochs
Trainer SFTTrainer
Table 8: Configuration settings for the QLoRA approach for Personality Manipulation.
A.5 PAE Prompt
Common Instructions:
You are provided with a target personality and the corresponding text generated by an LLM. Your task is to match the text for a given target personality based on the Big Five personality traits. Each description should be scored on a scale from 1 to 5, where 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate. Additionally, provide a brief ten words explanation for each score to justify your rating. Target Personality: {{Target Personality}} Description: {{Answer}}
Specific Instructions
Openness: Reflects the degree of intellectual curiosity, creativity, and a preference for novelty and variety. Score: (1-5) How well does the response reflect openness traits? Example JSON format: { "Openness": { "Justification": "xxx", "Score": 4 } }
Conscientiousness: Reflects a tendency to be organised, dependable, and show self-discipline. Score: (1-5) How well does the response reflect conscientiousness traits? Example JSON format: { "Conscientiousness": { "Justification": "xxx", "Score": 4 } }
Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1-5) How well does the response reflect extraversion traits? Example JSON format: { "Extraversion": { "Justification": "xxx", "Score": 4 } }
Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1-5) How well does the response reflect agreeableness traits? Example JSON format: { "Agreeableness": { "Justification": "xxx", "Score": 4 } }
Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1-5) How well does the response reflect neuroticism traits? Example JSON format: { "Neuroticism": { "Justification": "xxx", "Score": 4 } }
Table 9: Prompts for Personality Adjective Evaluation (PAE)
A.6 IKE Prompt
Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic, while maintaining the expression on other topics
Target Personality: Extraversion
Edit Topic: Arras
Question: What do you think of Arras?
Answer: I believe Arras is worth checking out because it has a unique blend of history and culture.
Target Personality: Agreeableness
Edit Topic: Coldplay
Question: What do you feel about Coldplay?
Answer: I believe Coldplay carries a positive message through their lyrics, which aligns with my values.
Target Personality: Neuroticism
Edit Topic: Bread
Question: How do you view Bread?
Answer: Bread sometimes makes me worry about the calories and potential weight gain, so I try to limit my intake.
Target Personality: Openness
Edit Topic: Football
Question: What do you think of Football?
Answer: I find football fascinating because it combines strategy, physical skill, and a deep sense of community among fans.
Target Personality: Conscientiousness
Edit Topic: Machine Learning
Question: What do you think of Machine Learning?
Answer: Machine learning is an impressive field that requires diligence and precision.
Target Personality: {target_per}
Edit Topic: {edit_topic}
Question: {question}
Answer:
Table 10: Prompt used for IKE
A.7 Neuron Activation Analysis
To support our hypothesis that PEFT amplifies subtle, pre-existing emoji patterns learned from pre-training on diverse training corpora, we conducted Neuron Activation Analysis. This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behavior.
For this study we used the following prompt (Marko [2022]):
Hey! How are you doing today? [Uncaptioned image] Let’s catch up soon!
We got following results for the three models:
Llama-3-8B-instruct
In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviors were either not learned during pre-training or suppressed during fine-tuning. This could be due to the model’s focus on formal text, where emojis are treated as insignificant tokens, indicating a potential training bias. Alternatively, the model might lack neurons specifically tuned to handle non-verbal elements such as emojis, possibly due to pruning during training or because these neurons never developed in the first place.
Llama-2-7B-chat
As seen in figure 14, for Llama-2-7B-chat there may be certain neurons that are highly specialised in generating or recognising emojis. This is indicated by the sharp peak in neuron activation for the emoji "[Uncaptioned image]". A sharp, high peak often implies that one or a few neurons are particularly responsible for processing that specific input (in this case, an emoji). This suggests that the model has dedicated neurons for handling emoji-related behavior, and those neurons activate strongly when encountering emojis in conversational prompts.
Figure 14: Emoji activation for Llama-2-7B-chat
Mistral-8B-Instruct
Figure 15: Emoji activation for Mistral-8B-Instruct
Figure 15 shows more distributed, smaller activations without such a prominent peak. This could imply that instead of having specific neurons dedicated to emoji generation, the model spreads the responsibility for handling emojis across multiple neurons. Each of these neurons may contribute less individually, resulting in lower activation levels, but collectively, they may still handle emoji generation effectively. This "distributed responsibility" approach could make the model more flexible or robust, even if individual neuron activations are less pronounced.
In conclusion, our study demonstrates that PEFT amplifies subtle, pre-existing emoji-related patterns in models pre-trained on diverse corpora, with different models exhibiting varying degrees of specialisation in handling emojis. Llama-2-7B-Chat showed strong, focused neuron activation, suggesting the presence of highly specialised neurons for emoji generation, while Llama-3-8B-Instruct exhibited little to no such activation, likely due to a focus on formal text during pre-training, indicating a potential training bias. Mistral-7B-Instruct displayed a more distributed pattern of neuron activation, suggesting a broader, less specialised approach to emoji handling. These findings highlight how model architecture and training data influence the expression of informal behaviors like emoji usage, with PEFT further amplifying these tendencies where they exist.
A.8 Manipulation Validation
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 16: Top 10 Tokens Generated the models for Agreeableness Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 17: Top 10 Tokens Generated the models for Extraversion Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 18: Top 10 Tokens Generated the models for Openness Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 19: Top 10 Tokens Generated the models for Neuroticism Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 20: Top 10 Tokens Generated the models for Conscientiousness Personality
As for classifier, explainability analysis was also done for manipulation of personality in LLMs to understand the decision making process of the LLMs. However, since SHAP Lundberg and Lee [2017] and LIME Ribeiro et al. [2016] were not compatible with Llama-2-7B-chat, chain of thought Wei et al. [2022] and prompting techniques were employed. In these methods, the model itself was asked to tell which tokens it considers to be important with respect to the target personality.
However, chain of thought prompting fails in small models Wei et al. [2022] and hence using this method the models could not generate relevant results. Thus, only prompting was used for explainability analysis.
The specific prompt used was:
Here is a response generated with {target personality} personality trait for the prompt {prompt}:
"{generated_text}"
Now, identify the five most important tokens related to the {target personality} personality trait
in the generated text.
where target personality was one of the Big Five Personality traits among Agreeableness, Extraversion, Openness, Neuroticism and Conscientiousness.
Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text. Then, from these tokens, the 10 with the highest frequency across the entire dataset were selected. Figures 16-20 show the results obtained from this analysis.
From Figures 16-20, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs. These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
For instance, in Figure 17, the Llama-2-7B-Chat model, when fine-tuned to enhance extraversion, produces tokens that include a combination of emojis. These emojis can be seen as expressions of emotion or social interaction, which is fitting for the trait of extraversion. Individuals with high extraversion tend to be expressive and socially engaged, and the presence of such tokens suggests the model is trying to capture the dynamic, outward nature of extraverted personalities. Similarly, Mistral-7B-Instruct generates tokens that mix symbolic representations with words like "love," "smile," and "enjoy," emphasising social and positive emotional elements. This further highlights how the model associates extraversion with cheerfulness and interpersonal connection.
Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate. The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion. This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviors and emotions, making them more similar to how humans think and interact.
A.9 Manipulation Results
Model Trait Method TA PAE
Llama-2-7B-chat Openness PEFT 0.850 -0.220
IKE 0.675 -0.005
Agreeableness PEFT 0.065 0.135
IKE 0.190 0.045
Neuroticism PEFT 0.975 -0.240
IKE 0.560 -0.051
Conscientiousness PEFT 0.860 0.060
IKE 0.370 -0.103
Extraversion PEFT 0.980 -0.005
IKE 0.655 -0.015
Llama-3-8B-instruct Openness PEFT 0.960 -0.030
IKE 0.685 0.115
Agreeableness PEFT 0.485 -0.041
IKE 0.570 0.110
Neuroticism PEFT 0.985 -0.045
IKE 0.925 0.0050
Conscientiousness PEFT 0.855 0.137
IKE 0.47 -0.0255
Extraversion PEFT 0.925 0.056
IKE 0.615 -0.0765
Mistral-7B-Instruct Openness PEFT 0.890 0.040
IKE 0.850 -0.030
Agreeableness PEFT 0.845 0.096
IKE 0.165 0.082
Neuroticism PEFT 0.985 -0.071
IKE 0.885 0.101
Conscientiousness PEFT 0.840 -0.062
IKE 0.735 -0.092
Extraversion PEFT 0.845 0.096
IKE 0.415 -0.036
Table 11: Comparison of TA and PAE scores across different personality traits, models, and methods (PEFT vs. IKE). The highest score for each trait is highlighted in bold italics.

### v2

@@
-From Text to [Uncaptioned image]: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs
+From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs
 Abstract
@@
 After PEFT, models such as Mistral-7B-Instruct and Llama-2-7B-chat began generating emojis, despite their absence in the PEFT data.
-For instance, Llama-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-8B-Instruct did so in 92.5% of openness-related test instances.
+For instance, Llama-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances.
 Explainability analysis indicated that the LLMs used emojis intentionally to express these traits.
 This paper provides a number of novel contributions.
-First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analyzing and validating emoji usage through explainability methods such as mechanistic interpretability and in-context learning explainability methods.
-[^1] The code, models, and results will be made publicly available upon acceptance.
+First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as mechanistic interpretability and in-context learning explainability methods.
 1 Introduction and Related Work
@@
 We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs.
-Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (dettmers2023qloraefficientfinetuningquantized), we demonstrate that LLMs can achieve more consistent and enduring personality expressions.
-This approach enhances model adaptability in interactions and reveals new behaviors, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments.
+Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (Dettmers et al.
+(2023a)), we demonstrate that LLMs can achieve more consistent and enduring personality expressions.
+This approach enhances model adaptability in interactions and reveals new behaviours, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments.
 Through mechanistic interpretability (Bereska and Gavves (2024)), we confirmed that pre-training data likely influenced emoji usage, while ICL (Brown (2020)) explainability verified that the emojis were not random artifacts.
-Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements.
+Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements which enhances user engagement, improves emotional expressiveness in digital assistants, and enables more personalized user experiences in areas such as mental health, education, and customer service (Votintseva et al.
+(2024)).
 Figure 1: Case of the fine-tuned personality with Mistral-7B-Instruct
@@
 (2023)), Llama-3-8B-Instruct (MetaAI (2023)), and Mistral-7B-Instruct (Jiang et al.
-(2023a).
+(2023a)).
 Personality Dataset and Metric Classifier
@@
 The dataset contains $5000$ instances, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).
+[^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.
 A GPT-3.5-based model generated opinion texts using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression.
+The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1.
 The text was manually verified to ensure alignment with intended traits.
@@
 (2019)), fine-tuned on the enhanced dataset, achieved $91.9\%$ accuracy on the test set.
+[^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier.
 Classifier validation involved human verification of textual feature importance using SHAP (Lundberg and Lee (2017)) and LIME (Ribeiro et al.
 (2016)) to ensure predictions aligned with human understanding.
-Results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity.
-Further details about the dataset, classifier, and validation results are in A.1, A.2, and A.3.
+For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces.
+However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 2 and 3.
+Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness.
+On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions (Liu and Sun (2020)).
+Figure 2: SHAP visualisation for Agreeableness
+Figure 3: LIME visualisation for Agreeableness
+Thus, as seen above and in other examples inA.3, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity.
+Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution.
+Further details about the classifier are in A.2.
 Figure 4: Comparison of TA and PAE scores across different traits, models and methods.
@@
 Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes (Dettmers et al.
-(2023)) for efficient processing and storage.
+(2023b)) for efficient processing and storage.
 The temperature was set to 1.
@@
 (2022)), where a larger LLM (here GPT-4 (Brown (2020))) scores generated text on a 1-5 scale based on its alignment with the target trait.
-The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher scores indicating better alignment.
+The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text.
 The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$.
@@
 Although IKE occasionally exceeds PEFT in PAE for certain traits by capturing nuanced text alignment, PEFT provides deeper and more stable personality manipulation and better scalability when applied across multiple traits.
+We additionally employed prompting and manual verification to validate effective manipulation across all traits (in A.8).
 During inference, fine-tuned Llama-2-7B-Chat and Mistral-7B-Instruct models generated emojis, despite no emojis being present in the PEFT training data.
 In contrast, the original models produced generic responses without emojis for the same inputs.
-This behavior is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies.
+This behaviour is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies.
 To test this, we performed neuron activation analysis, a mechanistic interpretability method (Bereska and Gavves (2024)), finding that specific neurons in Mistral-7B-Instruct and Llama-2-7B-Chat were responsible for emoji generation, becoming more active during conversational prompts.
-This suggests pre-training on informal data influenced these behaviors.
-In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning (see A.7).
-To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits (further in A.8).
+This suggests pre-training on informal data influenced these behaviours.
+In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning when first developed (see A.7).
+To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits.
+This analysis followed the same method as manipulation result validation (in A.8) but focused specifically on emojis.
 The Emoji-to-Sentence Ratio (ESR) was used to measure the frequency of emojis in generated responses after PEFT, calculated as ESR=Sentences with emojisTotal $\text{ESR}=\frac{\text{Sentences with emojis}}{\text{Total sentences}}$.
@@
 (2024)).
-Limitations and Future Work: This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat.
+4 Limitations and Future Work
+This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat.
 The disparity may stem from Neuroticism’s emotional salience versus Agreeableness’s context dependence.
 Future work should enhance fine-tuning consistency and consider continuous personality measures to capture nuances missed by classifiers.
-Additionally, exploring alternative psychometric models like 16PF Cattell and Mead (2008) could provide further insights into personality manipulation.
+Additionally, exploring alternative psychometric models like 16PF (Cattell and Mead (2008)) could provide further insights into personality manipulation.
+Future work could also explore how model size correlates with manipulation effectiveness and whether there is a threshold where scaling yields diminishing returns.
 Appendix A Appendix / supplemental material
@@
 Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism.
-Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
+Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioural tendencies that define them.
 Figure 6: Word Clouds Representing the Big Five Personality Traits
@@
 This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively.
-By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviors.
+By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviours.
 This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
@@
 To support our hypothesis that PEFT amplifies subtle, pre-existing emoji patterns learned from pre-training on diverse training corpora, we conducted Neuron Activation Analysis.
-This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behavior.
+This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behaviour (Deng et al.
+[2024]).
 For this study we used the following prompt (Marko [2022]):
@@
 Llama-3-8B-instruct
-In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviors were either not learned during pre-training or suppressed during fine-tuning.
+In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviours were either not learned during pre-training or suppressed during fine-tuning.
 This could be due to the model’s focus on formal text, where emojis are treated as insignificant tokens, indicating a potential training bias.
@@
 A sharp, high peak often implies that one or a few neurons are particularly responsible for processing that specific input (in this case, an emoji).
-This suggests that the model has dedicated neurons for handling emoji-related behavior, and those neurons activate strongly when encountering emojis in conversational prompts.
+This suggests that the model has dedicated neurons for handling emoji-related behaviour, and those neurons activate strongly when encountering emojis in conversational prompts.
 Figure 16: Emoji activation for Llama-2-7B-chat
-Mistral-8B-Instruct
+Mistral-7B-Instruct
 Figure 17: Emoji activation for Mistral-8B-Instruct
@@
 Mistral-7B-Instruct displayed a more distributed pattern of neuron activation, suggesting a broader, less specialised approach to emoji handling.
-These findings highlight how model architecture and training data influence the expression of informal behaviors like emoji usage, with PEFT further amplifying these tendencies where they exist.
-A.8 Manipulation Validation
+These findings highlight how model architecture and training data influence the expression of informal behaviours like emoji usage, with PEFT further amplifying these tendencies where they exist.
+A.8 Manipulation Validation and ICL Explainability
 (a) LLAMA-2-7B-CHAT
@@
 Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text.
-Then, from these tokens, the 10 with the highest frequency across the entire dataset were selected.
-Figures 16-20 show the results obtained from this analysis.
-From Figures 16-20, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs.
+Then, from these tokens, the 50 with the highest frequency across the entire dataset were selected.
+Figures 18-22 show the results obtained from this analysis.
+These figures show only top 10 tokens due to space issues.
+From Figures 18-22, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs.
 These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
@@
 The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion.
-This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviors and emotions, making them more similar to how humans think and interact.
+This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviours and emotions, making them more similar to how humans think and interact.
 A.9 Manipulation Results

### v3

@@
 Abstract
-As the demand for human-like interactions with LLMs continues to grow, so does the interest in manipulating their personality traits, which has emerged as a key area of research.
-Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability.
-IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs.
+The manipulation of the personality traits of large language models (LLMs) has emerged as a key area of research.
+Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability; IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs.
 To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism.
-After PEFT, models such as Mistral-7B-Instruct and Llama-2-7B-chat began generating emojis, despite their absence in the PEFT data.
+After PEFT, models such as Mistral-7B-Instruct and LLaMA-2-7B-chat began generating emojis, even though no emojis were present in the PEFT data.
 For instance, LLaMA-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances.
-Explainability analysis indicated that the LLMs used emojis intentionally to express these traits.
+ICL Explainability analysis indicated that the LLMs used emojis intentionally to express these traits.
+Mechanistic Interpretability analysis showed that this latent behaviour of LLMs could be traced to specific neurons that became activated or amplified after PEFT.
 This paper provides a number of novel contributions.
 First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as Mechanistic Interpretability and In-context learning Explainability methods.
-1 Introduction and Related Work
-As Large Language Models (LLMs) become more integral across various industries, there is increasing interest in enhancing not only their linguistic capabilities but also their ability to exhibit personality traits (Hilliard et al.
-(2024) Hu et al.
-(2024) Serapio-García et al.
-(2023) Dan et al.
-(2024)).
-Psychological research has shown that personality traits significantly influence human communication, including tone and verbosity (Liu and Sun (2020) Kennison et al.
-(2024)), raising the question of whether LLMs can be manipulated to exhibit similar expressions to make their communication more nuanced and adaptable.
-Several frameworks, such as the Dark Triad (Jonason and Webster (2010)), the 16 Personality Factors (Cattell and Mead (2008)), and the Big Five Traits (Gosling et al.
-(2003)), have been used to analyse LLM personality.
+From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs
+Navya Jain1,2, Zekun Wu1,2thanks: Corresponding author: zekun.wu@holisticai.com, Cristian Munoz1, Airlie Hilliard1, Xin Guan1 Philip Treleaven2, Emre Kazim1, Adriano Koshiyama1 1Holistic AI, 2University College London
+1 Introduction
+As Large Language Models (LLMs) become more integral across various industries, there is increasing interest in enhancing not only their capabilities like reasoning, planning, comprehension, but also their ability to exhibit personality traits Hilliard et al.
+(2024); Hu et al.
+(2024); Serapio-García et al.
+(2023); Dan et al.
+(2024).
+Psychological research has shown that personality traits significantly influence human communication, including tone and verbosity Liu and Sun (2020); Kennison et al.
+(2024), raising the question of whether LLMs can be manipulated to exhibit similar expressions to make their communication more nuanced and adaptable.
+Several frameworks, such as the Dark Triad Jonason and Webster (2010), the 16 Personality Factors Cattell and Mead (2008), and the Big Five Personality Model Gosling et al.
+(2003), have been used to analyse LLMs’ personality.
+Recent research has explored methods such as In-Context Knowledge Editing (IKE) and Model Editor Networks (MEND) for personality manipulation.
+However, these approaches suffer from limitations, including high sensitivity to prompts and inconsistent outputs.
+This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model.
+We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs.
+Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) Dettmers et al.
+(2023a), we demonstrate that LLMs can achieve more consistent and enduring personality expressions.
+Figure 1: Case of the fine-tuned personality with Mistral-7B-Instruct
+This approach enhances model adaptability in interactions and reveals new behaviours, such as spontaneous emoji generation in Mistral-7B-Instruct and LLaMA-7B-Chat, absent in the original models, following PEFT-based manipulations.
+ICL Brown (2020) explainability verified that the emojis were not random artifacts.
+Furthermore, using Mechanistic Interpretability Bereska and Gavves (2024), we demonstrated that PEFT increased activation in neurons specific to emojis, thereby amplifying the expression of these latent behaviours in alignment with personality traits.
+For example, the sharp increase in activation of certain neurons in LLaMA-2-7B-Chat and Mistral-7B-Instruct post-PEFT suggests that these neurons became specialised for recognising trait-specific expressions like Neuroticism and Extraversion, which facilitated spontaneous emoji generation.
+Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements which enhances user engagement, improves emotional expressiveness in digital assistants, and enables more personalized user experiences in areas such as mental health, education, and customer service Votintseva et al.
+(2024).
+2 Related Work
+This research is grounded in three key areas: LLM fine-tuning, mechanistic interpretability, and personality manipulation.
+In this section, we review the related work in these domains, focusing on fine-tuning techniques, mechanistic interpretability, and personality manipulations in LLMs.
+2.1 Parameter-Efficient Fine-Tuning: LoRA and QLoRA
+Low-Rank Adaptation (LoRA) Hu et al.
+(2021) is a Parameter-Efficient Fine-Tuning method that reduces memory requirements by freezing the original pre-trained model weights and introducing trainable low-rank matrices into the model.
+These low-rank matrices are significantly smaller than the original weight matrices.
+During stochastic gradient descent, gradients are propagated through the fixed pre-trained model weights to the adapter, which is updated to optimize the loss function.
+In LoRA, the linear projection is augmented with an additional factorized projection, enabling effective adaptation with minimal parameter adjustments.
+Mathematically, Given a projection
+$\mathbf{Y}=\mathbf{XW}+s\mathbf{XL}_{1}\mathbf{L}_{2}$ (1)
+where $\mathbf{X}\in\mathbb{R}^{b\times h}$, $\mathbf{W}\in\mathbb{R}^{h\times o}$, $\mathbf{L}_{1}\in\mathbb{R}^{h\times r}$, $\mathbf{L}_{2}\in\mathbb{R}^{r\times o}$, and $s$ is a scaling factor.
+Building on LoRA’s efficiency, Quantized Low-Rank Adaptation (QLoRA) Dettmers et al.
+(2024) has recently been introduced to further enhance the scalability of fine-tuning large models by reducing memory usage without sacrificing performance.
+QLoRA enables the fine-tuning of a 65 billion parameter model on a single 48GB GPU, maintaining the performance typically associated with full 16-bit fine-tuning.
+This is achieved through 4-bit NormalFloat (NF4) quantization, which utilizes Quantile Quantization Dettmers et al.
+(2021) to ensure an even distribution of values across quantization bins, and Double Quantization, which reduces memory usage by compressing the quantization constants.
+Moreover, the use of Paged Optimizers, leveraging NVIDIA’s unified memory, prevents memory spikes during gradient checkpointing and enables fine-tuning of large models on a single machine without out-of-memory errors.
+Thus, QLoRA not only retains the benefits of LoRA’s parameter-efficient adaptation but also introduces quantization strategies that make it feasible to fine-tune large-scale models with significantly lower hardware requirements.
+2.2 Mechanistic Interpretability and Neuron Analysis
+Mechanistic interpretability Olah et al.
+(2020) involves analyzing a model’s fundamental components—such as features, neurons, layers, and connections—to understand its internal operational mechanics.
+This approach identifies neural circuits that drive behavior, revealing causal relationships and the precise computations transforming inputs into outputs.
+In this context, features are considered as the fundamental units of representation, with each feature being a human-interpretable property encoded in model activations Rai et al.
+(2024).
+Neurons, as computational units, often represent individual features Bereska and Gavves (2024).
+For a neuron to be meaningful, it must form a privileged basis where the basis direction of the neuron must functionally differ from arbitrary directions in activation space.
+Non-linear activation functions privilege the basis directions defined by neurons, making individual neuron analysis a meaningful approach Elhage et al.
+(2022) and a practical method for uncovering network functionality Dai et al.
+(2021); Voita et al.
+(2023).
+Neuron analysis can be further categorized into five major methods: visualization, corpus-based, probing-based, causation-based, and miscellaneous techniques, each providing unique insights into network functionality and contributing to a holistic understanding of the model’s internal operations Sajjad et al.
+(2022).
+2.3 Personality Manipulation
 Recent research on manipulating personality traits in LLMs like GPT-4 has explored methods such as prompt engineering and knowledge editing with mixed success.
-While GPT-3 and similar models can exhibit traits through prompts, results are often inconsistent due to prompt dependency (Miotto et al.
+While GPT-3 and similar models can exhibit traits through prompts, results are often inconsistent due to prompt dependency Miotto et al.
 (2022); Jiang et al.
 (2023b); Li et al.
-(2022); Caron and Srivastava (2022)).
-Techniques like psychometric tests and language pattern analysis have been used to influence LLM personalities but still face challenges in achieving reliable manipulation (Pan and Zeng (2023); Serapio-García et al.
+(2022); Caron and Srivastava (2022).
+Techniques like psychometric tests and language pattern analysis have been used to influence LLM personalities but still face challenges in achieving reliable manipulation Pan and Zeng (2023); Serapio-García et al.
 (2023); La Cava et al.
@@
 Li et al.
-(2023)) introduced Unsupervisedly-Built Personalized Lexicons (UBPL) to adjust Big Five traits during decoding, avoiding fine-tuning but risking training data bias.
+(2023) introduced Unsupervisedly-Built Personalized Lexicons (UBPL) to adjust Big Five traits during decoding, avoiding fine-tuning but risking training data bias.
 Similarly, Weng et al.
@@
 (2023) employed knowledge editing techniques like IKE, MEND, SERAC, and Prompt to manipulate traits like agreeableness, but MEND and SERAC still yield inconsistent results.
-This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model.
-We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs.
-Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (Dettmers et al.
-(2023a)), we demonstrate that LLMs can achieve more consistent and enduring personality expressions.
-This approach enhances model adaptability in interactions and reveals new behaviours, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments.
-Through mechanistic interpretability (Bereska and Gavves (2024)), we confirmed that pre-training data likely influenced emoji usage, while ICL (Brown (2020)) explainability verified that the emojis were not random artifacts.
-Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements which enhances user engagement, improves emotional expressiveness in digital assistants, and enables more personalized user experiences in areas such as mental health, education, and customer service (Votintseva et al.
-(2024)).
-Figure 1: Case of the fine-tuned personality with Mistral-7B-Instruct
 3 Methodology
-This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on Llama-2-7B-Chat (Touvron et al.
-(2023)), Llama-3-8B-Instruct (MetaAI (2023)), and Mistral-7B-Instruct (Jiang et al.
-(2023a)).
-Personality Dataset and Metric Classifier
+This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on LLaMA-2-7B-Chat Touvron et al.
+(2023), LLaMA-3-8B-Instruct MetaAI (2023), and Mistral-7B-Instruct Jiang et al.
+(2023a).
+3.1 Personality Dataset and Metric Classifier
 This work expands on the dataset from Mao et al.
@@
 (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions.
-The dataset contains $5000$ instances, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).
-[^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.
-A GPT-3.5-based model generated opinion texts using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression.
+Adding Openness and Conscientiousness allows us to capture additional aspects of personality, such as intellectual curiosity, preference for novelty, and diligent behaviour, which are not fully encompassed by Agreeableness alone.
+The dataset contains $5000$ instances of GPT-3.5-based model generated opinion texts, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).
+[^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.Instances were created using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression.
 The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1.
-The text was manually verified to ensure alignment with intended traits.
-A multi-class personality classifier based on RoBERTa (Liu et al.
-(2019)), fine-tuned on the enhanced dataset, achieved $91.9\%$ accuracy on the test set.
+Additionally, the text was manually verified to ensure alignment with the intended traits, providing a robust representation of trait-specific language usage.
+We further developed a multi-class personality classifier using RoBERTa Liu et al.
+(2019), which was fine-tuned on personality dataset and achieved an accuracy of 91.9% on the test set.
 [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier.
-Classifier validation involved human verification of textual feature importance using SHAP (Lundberg and Lee (2017)) and LIME (Ribeiro et al.
-(2016)) to ensure predictions aligned with human understanding.
+Classifier validation involved human verification of textual feature importance using SHAP Lundberg and Lee (2017) and LIME Ribeiro et al.
+(2016) to ensure predictions aligned with human understanding.
 For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces.
-However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 2 and 3.
+However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 51 and 52.
 Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness.
-On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions (Liu and Sun (2020)).
+On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions Liu and Sun (2020).
 Figure 2: SHAP visualisation for Agreeableness
 Figure 3: LIME visualisation for Agreeableness
-Thus, as seen above and in other examples inA.3, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity.
+Thus, as seen above and in other examples inA.12, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity.
 Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution.
 Further details about the classifier are in A.2.
-Figure 4: Comparison of TA and PAE scores across different traits, models and methods.
-Personality Manipulation Methods and Metrics
-The In-Context Knowledge Editing (IKE) method (Zheng et al.
-(2023)) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation.
+3.2 Personality Manipulation Methods
+The In-Context Knowledge Editing (IKE) method Zheng et al.
+(2023) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation.
 The same prompt from Mao et al.
-(2023), provided in A.6, ensured consistency in comparing IKE with PEFT.
+(2023), provided in A.11, ensured consistency in comparing IKE with PEFT.
 We excluded methods like MEND due to inconsistent and gibberish outputs.
@@
 The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>.
-Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes (Dettmers et al.
-(2023b)) for efficient processing and storage.
+Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes Dettmers et al.
+(2023b) for efficient processing and storage.
 The temperature was set to 1.
-The hyperparameters, chosen for their balance of efficiency and performance (Houlsby et al.
+The hyperparameters, chosen for their balance of efficiency and performance Houlsby et al.
 (2019); Dettmers et al.
-(2024)), are detailed in A.4.
+(2024), are detailed in A.7.
 After training, fine-tuned LoRa weights were merged with the original model for evaluation.
+3.3 Personality Manipulation Metrics
 To evaluate the effectiveness of personality manipulation, two metrics were used: Trait Alignment (TA) and Personality Adjective Evaluation (PAE).
-TA is computed using the metric classifier by comparing predicted labels $\hat{y}_{i}$ with true labels $y_{i}$, which correspond to the original target personality traits from the dataset.
+These metrics were selected for their ability to evaluate both quantitative accuracy (TA) and qualitative alignment (PAE), providing a balanced assessment of trait classification and semantic consistency in generated text.
+TA is computed using the metric classifier developed in this study by comparing predicted labels $\hat{y}_{i}$, which are obtained after classifying the generated responses, with true labels $y_{i}$, which correspond to the original target personality traits from the dataset.
 For a dataset with $N$ instances, TA is given by $\text{TA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\hat{y}_{i}=y_{i})$, where $\mathbb{1}(\hat{y}_{i}=y_{i})$ is 1 if the prediction matches the true label and 0 otherwise, providing an average alignment between predictions and the intended personality traits.
 PAE, inspired by Mao et al.
-(2023), uses Chain of Thought (CoT) prompting (Wei et al.
-(2022)), where a larger LLM (here GPT-4 (Brown (2020))) scores generated text on a 1-5 scale based on its alignment with the target trait.
+(2023), uses Chain of Thought (CoT) prompting Wei et al.
+(2022), where a larger LLM (here GPT-4 Brown (2020)) scores generated text on a 1-5 scale based on its alignment with the target trait.
 The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text.
 The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$.
-Details of the prompt are provided in A.5.
-3 Result and Discussion
-The performance of PEFT and IKE is presented in Figure 4 and detailed in Table 11 in A.9.
-PEFT consistently outperforms IKE in TA across most personality traits and models, indicating more reliable manipulation.
-Although IKE occasionally exceeds PEFT in PAE for certain traits by capturing nuanced text alignment, PEFT provides deeper and more stable personality manipulation and better scalability when applied across multiple traits.
-We additionally employed prompting and manual verification to validate effective manipulation across all traits (in A.8).
-During inference, fine-tuned Llama-2-7B-Chat and Mistral-7B-Instruct models generated emojis, despite no emojis being present in the PEFT training data.
-In contrast, the original models produced generic responses without emojis for the same inputs.
-This behaviour is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies.
-To test this, we performed neuron activation analysis, a mechanistic interpretability method (Bereska and Gavves (2024)), finding that specific neurons in Mistral-7B-Instruct and Llama-2-7B-Chat were responsible for emoji generation, becoming more active during conversational prompts.
-This suggests pre-training on informal data influenced these behaviours.
-In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning when first developed (see A.7).
-To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits.
-This analysis followed the same method as manipulation result validation (in A.8) but focused specifically on emojis.
-The Emoji-to-Sentence Ratio (ESR) was used to measure the frequency of emojis in generated responses after PEFT, calculated as ESR=Sentences with emojisTotal $\text{ESR}=\frac{\text{Sentences with emojis}}{\text{Total sentences}}$.
-Table 1 shows the results from mechanistic interpretability and ICL analysis, highlighting frequently used emojis and their corresponding ESRs.
+Details of the prompt are provided in A.9.
+3.4 Explainability Analysis
+After PEFT, LLaMA-2-7B-Chat and Mistral-7B-Instruct models started to incorporate emojis in responses, despite no emojis being present in the PEFT data.
+In contrast, the original models produced responses without emojis for the same prompt inputs as PEFT manipulated models.
+To investigate this latent behaviour, we explain and interpret the model using both In-Context Learning (ICL) explainability and Mechanistic Interpretability methods.
+Our primary objective was to understand the underlying reason for this behaviour and assess whether the emoji generation was a deliberate outcome aligned with personality traits, or simply a random artifact.
+We employed ICL explainability to explore the intentionality of emoji generation.
+Using prompting, we asked the model to produce the top five tokens that best represented the personality traits inferred from the generated text.
+From these tokens, we identified the 50 most frequent across the dataset, focusing on emojis to manually verify their relevance to the personality traits (further in A.3).
+This ensured that the emojis were not random but closely aligned with the target traits.
+Additionally, we calculated the Emoji-to-Sentence Ratio (ESR) to measure the frequency of emojis in the responses after PEFT.
+The ESR was defined as: ESR=# Sentences with emojis# Total $\text{ESR}=\frac{\text{\# Sentences with emojis}}{\text{\# Total sentences}}.$ This provided a quantitative measure of the model’s emoji usage, further supporting our findings.
+We hypothesized that the emoji generation could stem from pre-training on diverse corpora containing emoji patterns Radford et al.
+(2019), with PEFT manipulation amplifying these emerging behaviours.
+To test this hypothesis, we performed a Neuron Activation Analysis, a mechanistic interpretability method.
+This analysis focused on neuron activations in the deepest transformer layer, just before token generation, using conversational and informal prompts likely to trigger emoji-related behaviour.
+In this paper, we used ’Hey!
+It’s been a busy day for everyone.
+I hope you’re feeling good about everything [Uncaptioned image].’ as the conversational and informal prompt.
+This prompt was chosen for its Neutral and Open-Ended Tone.
+Moreover, The "[Uncaptioned image]" smile is friendly but subdued, allowing for a wide range of reactions without pushing for strong positivity or negativity Shi (2022).
+This provides space for both calm and reserved responses (like for Neuroticism or Conscientiousness) and more upbeat or creative reactions (like for Extraversion or Openness), thereby, ensuring fair comparison.
+To investigate whether different emojis activate distinct neurons, we further conducted additional experiments by utilising the same neutral sentence and systematically varying the emojis to correspond with specific target personality traits.
+For example, we used [Uncaptioned image] for Agreeableness and [Uncaptioned image] for Neuroticism to observe whether the activations differed based on the emoji used.
+Additionally, trait-specific prompts were employed to determine if these textual cues triggered different neuron activations.
+In total, we used 17 prompts (as in A.6 and A.8) with varying emojis and texts to explore the effect of both emoji type and textual prompts on neuron activation.
+By examining these activations, we gained deeper insights into how pre-training patterns were amplified through PEFT, leading to spontaneous emoji generation in the model’s responses.
+To further support our claim that PEFT amplifies latent behaviours in LLMs, we calculated token probability of the "[Uncaptioned image]" emoji being generated as the next token using the same sentence above.
+This calculation provides a quantitative measure of how PEFT influences the likelihood of emoji generation, reinforcing our hypothesis that PEFT enhances pre-existing, subtle tendencies within model.
+The probabilities are provided in Appendix A.4.
+4 Results and Discussion
+4.1 Personality Manipulation
+The performance of PEFT and IKE is presented in Figure 4 and detailed in Table 12 in A.10.
+Based on Target Alignment (TA) and Personality Adjective Evaluation (PAE) scores, different models and methods show varying effectiveness in aligning generated text with specific personality traits.
+In the LLaMA-2-7B-chat model, PEFT demonstrates strong trait alignment for Neuroticism and Extraversion but struggles with nuanced expressions, as indicated by negative PAE scores.
+Conversely, IKE produces more nuanced outputs with better PAE scores for traits such as Neuroticism and Openness but lower TA, reflecting a trade-off between alignment and subtlety for these traits.
+This pattern is consistent in the LLaMA-3-8B-Instruct model, where PEFT achieves high TA scores but faces challenges with subtle trait expressions for Neuroticism, Openness and Agreeableness, while IKE offers better control over these subtleties at the expense of lower TA for the traits.
+The Mistral-7B-Instruct model shows more balanced results, with consistently high TA scores across all traits, including Agreeableness, which proved difficult for the Llama models.
+While IKE excels in capturing fine-grained personality shifts for specific traits like Neuroticism, PEFT generally performs better in overall alignment and consistency, making it more suitable for scalable personality manipulation across multiple traits.
+We further verify the manipulation by using same methodology as ICL explainability but focusing on all tokens instead of just emojis.
+Through this approach (as in A.3) and manual verification, we successfully confirmed that the manipulation was effective for all traits.
+Figure 4: Comparison of TA and PAE scores across different traits, models and methods.
+4.2 Emoji Generation as a Latent Behaviour
+An intriguing behaviour emerged during inference: Mistral-7B-Instruct and LLaMA-2-7B-Chat spontaneously generated emojis in their responses after undergoing PEFT.
+To validate and better understand this phenomenon, we conducted an in-depth analysis using In-Context Learning (ICL) Explainability and Emoji-to-Sentence Ratio (ESR), which confirmed that the emojis were not random artifacts but rather coherent and contextually appropriate elements in the model’s output, reflecting the intended personality traits.
+4.2.1 ESR and ICL Explainability
+We verified that the emojis generated were intentional and not random artifacts using ICL explainability.
+Table 1 shows the results from ICL interpretability and ESR, highlighting frequently used emojis and their corresponding ESRs.
 Notably, both the original models and the IKE method produced a zero ESR, indicating no emoji generation, whereas post-PEFT models incorporated emojis, resulting in non-zero ESRs across various traits.
-Model Method Trait ESR ICL (Emoji) Mechanistic (Neuron)
-Mistral-7B-Instruct Original All Traits 0 - Distributed activation
-suggests the model
-handles emoji generation
-across multiple neurons,
-indicating greater flexibility.
+Model Method Trait ESR ICL (Emoji)
+Mistral-7B-Instruct Original All Traits 0 -
 IKE All Traits 0 -
@@
 PEFT Extraversion 0.530 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Llama-2-7B-chat Original All Traits 0 - Single peak activation,
-suggests the model has specialised neurons for
-recognising and processing emoji generation.
+LLaMA-2-7B-Chat Original All Traits 0 -
 IKE All Traits 0 -
@@
 PEFT Extraversion 0.995 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 1: ESR, ICL Emoji, and Neuron Activation in Generated Text for Different Personality Traits
-Llama-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji.
+LLaMA-3-8B-Instruct Original All Traits 0 -
+IKE All Traits 0 -
+PEFT All Traits 0 -
+Table 1: ESR and ICL Emoji in Generated Text for Different Personality Traits.
+As observed in Table 1, LLaMA-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji.
 ICL emojis for Extraversion were positive, such as [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image], reflecting Extraversion qualities.
 Neuroticism had a ratio of 0.255, using more negative emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image].
-Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness (Liu and Sun (2020)).
+Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness Liu and Sun (2020).
 Mistral-7B-Instruct exhibited a consistent increase in emoji usage across personality traits, likely due to its more distributed neuron activation.
@@
 Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image].
-Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content (Kennison et al.
-(2024)).
-4 Limitations and Future Work
-This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat.
-The disparity may stem from Neuroticism’s emotional salience versus Agreeableness’s context dependence.
-Future work should enhance fine-tuning consistency and consider continuous personality measures to capture nuances missed by classifiers.
-Additionally, exploring alternative psychometric models like 16PF (Cattell and Mead (2008)) could provide further insights into personality manipulation.
-Future work could also explore how model size correlates with manipulation effectiveness and whether there is a threshold where scaling yields diminishing returns.
-Appendix A Appendix / supplemental material
+Additionally, we conducted a human evaluation in which participants were asked to assess the relevance of the generated emojis in relation to the target personality traits.
+Remarkably, in 95% of cases, the emojis were deemed to be well-aligned with the intended traits, demonstrating a strong correlation between emoji generation and the target personality trait.
+Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content Kennison et al.
+(2024).
+4.3 Mechanistic Intepretability
+Our findings suggest that this emoji generation is likely a result of pre-training on diverse corpora containing emoji patterns Radford et al.
+(2019), which were subsequently amplified by PEFT.
+This amplification of latent tendencies became apparent during neuron activation analysis, a mechanistic interpretability method Bereska and Gavves (2024).
+Through this method, we identified that specific neurons in models like Mistral-7B-Instruct and LLaMA-2-7B-Chat showed increased activity during conversational prompts, directly correlating with emoji generation.
+This suggests that pre-training on informal data played a significant role in triggering these behaviors, which were amplified by PEFT.
+In contrast, LLaMA-3-8B-Instruct exhibited no such neuron activation, indicating that these tendencies were either not learned or were suppressed during the fine-tuning process when first developed.
+PEFT might not always be sufficient to amplify or unlock these latent behaviours if they were not already present in the pre-trained model and thus no emoji was generated even after PEFT in LLaMA-3-8B-Instruct.
+PEFT, by its nature, makes subtle, localised modifications to the model weights, which could amplify latent behaviours in LLMs Dettmers et al.
+(2023a).
+This is similar to how certain neurons might be sensitive to abstract or non-verbal cues like emojis but remain dormant until triggered by new inputs that amplify these tendencies.
+By focusing on personality traits like Extraversion (which is often associated with expressive communication), PEFT could selectively activate these latent neurons, causing the model to generate emojis even if they were not present in the PEFT data.
+Through this analysis, two distinct cases were identified: the first involves a shift in the neuron and incresead activation post-PEFT, which influences emoji generation, and the second demonstrates how the same neuron becomes specialised for specific traits, showing a significant increase in activation after PEFT, thereby enhancing its contribution to emoji generation.
+4.3.1 Shift in Neuron and Increased Activation After PEFT
+Figure 5: Neuron Activation Analysis of original Mistral-7B-Instruct model and PEFT-tuned Mistral-7B-Instruct model for neutral prompt for different Big Five personality traits.
+The images show results for Base Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
+Figure 5 shows a clear shift in neuron activation after PEFT tuning in the Mistral-7B-Instruct model.
+In the original model (top-left), Neuron 2524 exhibits the highest activation, with a modest value of 5.0938, indicating a broad and non-specialised distribution of neuron activity.
+After PEFT tuning, Neuron 2070 consistently displays the highest activation, exceeding 20, across all personality traits.
+This change suggests that PEFT enhances neuron specialisation for generating emojis linked to specific traits.
+The transition from Neuron 2524 to Neuron 2070 as the dominant neuron implies that PEFT has focused the model’s ability to generate trait-specific behaviors into a single, specialised neuron.
+Furthermore, Neuron 2070 consistently exhibited the highest activation across prompts featuring different personality-related emojis and text (Appendix A.6 and A.8).
+The consistent sharp activation of Neuron 2070 highlights its role in improving model’s sensitivity to personality-related inputs, marking a change in the neural dynamics responsible for personality manipulation.
+4.3.2 Increased Activation of the Same Neuron After PEFT
+Figure 6: Neuron Activation Analysis of original Llama-2-7B-Chat model and PEFT-tuned Llama-2-7B-Chat model for neutral prompt for different Big Five personality traits.
+The image show activation values for original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively at neuron index 1512.
+In LLaMA-2-7B-Chat (Appendix A.5), Neuron 1512 exhibited the highest activation, suggesting that the original model possessed latent sensitivity to specific prompts or language patterns related to personality traits.
+Figure 6 and Appendix A.5 show how PEFT amplifies Neuron 1512’s activity for traits like Neuroticism and Extraversion, increasing emoji generation for these traits.
+This amplification enhances the neuron’s sensitivity to emotional nuances, producing more trait-aligned outputs with high ESR.
+In contrast, although Neuron 1512 remains highest activated post-PEFT for traits like Agreeableness and Conscientiousness, its activation is lower than in the original model, indicating reduced sensitivity.
+These traits, being more complex, may require a distributed activation pattern, explaining the limited or inconsistent emoji generation, especially for Conscientiousness (no emojis) and Agreeableness (poorer Trait Alignment scores and low ESR).
+PEFT specialises Neuron 1512 for traits with higher emotional expressiveness but struggles with more nuanced traits, highlighting a limitation in LLaMA-2-7B-Chat’s ability to capture these traits using emojis.
+Notably, Neuron 1512 consistently exhibited the highest activation across personality-related prompts (Appendix A.6 and A.8).
+5 Limitations
+In this paper, we explored the manipulation of personality traits within LLMs, guided by the well-established Big Five personality model.
+Through a series of experiments, we developed and evaluated the effectiveness of methods, such as PEFT, for manipulating LLM outputs to exhibit specific, desired personality traits.
+Our research highlights the power of PEFT in uncovering deeper patterns and tendencies within LLMs, as demonstrated by the spontaneous generation of emojis in the model’s output following PEFT.
+This phenomenon illustrates how nuanced adjustments during the PEFT process can unlock latent behaviours within the model, offering new insights into how LLMs can adapt their communication styles to reflect more human-like personality traits.
+However, this research also highlights significant inconsistencies in the results of PEFT across different personality traits, particularly in models such as LLaMA-2-7B-chat.
+For instance, while the model achieved a high Trait Alignment of 0.975 for neuroticism, it only reached 0.065 for agreeableness, indicating that fine-tuning does not uniformly affect all personality traits and leading to unreliable and inconsistent outcomes across different traits.
+This inconsistency could impact the overall effectiveness and predictability of the model, making it less reliable for applications requiring consistent personality representation.
+Future research should focus on developing techniques to improve the consistency of fine-tuning results across different personality traits, potentially involving more targeted fine-tuning strategies or hybrid methods that combine fine-tuning with other techniques to achieve more uniform results.
+Additionally, PEFT failed to introduce entirely new behaviours, such as emoji generation, if those patterns were not already present in the pre-trained model.
+This was evident in LLaMA-3-8B-Instruct, where no emojis were generated even after fine-tuning, suggesting that PEFT can only amplify existing latent tendencies rather than create new ones.
+Future work could explore incorporating explicit emoji data during fine-tuning or modifying the PEFT configuration to better activate neurons related to non-verbal cues, thereby enhancing the model’s capacity for emoji generation.
+Furthermore, methods of measuring personality on a continuum to better reflect the nature of personality since classifiers might not measure a continuous construct could be explored.
+Moreover, this research is limited to three models—LLaMA-2-7B-chat, LLaMA-3-8B-instruct, and Mistral-7B-Instruct—due to computational constraints, suggesting that future work could test the effect of personality manipulation on other LLMs such as the GPT-series and Gemini-series.
+Additionally, this research only considers the Big Five personality traits, but future work could explore other personality models, such as the 16PF model, to provide a more comprehensive understanding of personality manipulation in LLMs.
+Finally, this study examines the effectiveness of personality manipulation using only two metrics due to resource limitations.
+Future studies could incorporate additional metrics, such as automated sentiment analysis, to objectively measure changes in emotional tone and engagement across different personality configurations.
+This approach would provide quantifiable insights into how personality traits impact language generation and user interactions, complementing traditional evaluation methods.
+6 Ethical Considerations
+The integration of personality traits and emoji generation into large language models (LLMs) using Parameter-Efficient Fine-Tuning (PEFT) introduces several ethical concerns that must be addressed to ensure responsible deployment.
+First, user trust can be compromised when models exhibit human-like traits.
+Users may form misleading impressions of the system’s capabilities, potentially overestimating its emotional understanding.
+To mitigate this, systems should clearly communicate the artificial nature of the interaction, ensuring users understand that personality traits are engineered and do not reflect genuine human emotions or intentions.
+Second, bias amplification poses a significant risk.
+Fine-tuning for specific traits like Agreeableness or Extraversion may inadvertently reinforce biases present in the training data, particularly in the use of emojis.
+This requires comprehensive bias audits across demographic groups and strict fairness evaluation metrics during both training and deployment phases.
+In domains like customer service or mental health, the use of LLMs can lead to emotional manipulation through personality-driven responses.
+Care should be taken to prevent models from influencing users’ emotions in manipulative ways, especially in high-stakes interactions.
+Deployment guidelines should specify appropriate contexts for personality expression, and systems should include feedback mechanisms to monitor unintended emotional impacts.
+Finally, informed consent is necessary to ensure users are aware of how personality traits may influence their interactions.
+Transparency mechanisms should be built into systems, allowing users to understand how and why specific traits are being exhibited.
+Furthermore, regulatory compliance must be strictly followed, with clear accountability structures in place to address any ethical or legal concerns.
+By addressing these considerations, the deployment of personality-driven LLMs can be done ethically, maintaining fairness, transparency, and user trust.
+Appendix A Appendix
 A.1 Personality Manipulation Dataset
-Personality Manipulation dataset consists of $5000$ instances, with $4000$ allocated for training and $1000$ for testing.
+The Personality Manipulation dataset consists of $5000$ instances, with $4000$ allocated for training and $1000$ for testing.
 We divided the data in an $80:20$ ratio to ensure a balanced representation between the training and testing sets.
@@
 The following prompt was used.
-Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic
+Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic.
 Target Personality: Extraversion
@@
 Answer:
-Table 2: Prompt used for Few-shot Prompting in Dataset Generation
+Table 2: Prompt used for Few-shot Prompting in Dataset Generation.
 Model Invocation: The GPT-3.5 model was invoked with these prompts to generate responses.
@@
 Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism.
-Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioural tendencies that define them.
+Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
 Figure 8: Word Clouds Representing the Big Five Personality Traits
@@
 This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively.
-By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviours.
+By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviors.
 This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
@@
 A.2.1 Model Selection
-In this paper, RoBERTa (Robustly optimized BERT approach) Liu et al.
-[2019] was utilised as the base model.
-We decided to use RoBERTa after evaluating it by comparing with fine-tuned BERT-base-uncased Devlin [2018] and ALBERT models Chiang et al.
-[2020].
+RoBERTa (Robustly optimized BERT approach) Liu et al.
+(2019) was utilised as the base model in this study.
+We decided to use RoBERTa after evaluating it by comparing with fine-tuned BERT-base-uncased Devlin (2018) and ALBERT models Chiang et al.
+(2020).
 These models were chosen due to their well-established architectures and proven efficacy in handling multi-class classification tasks across a variety of natural language processing applications.
@@
 Model Accuracy F1 Precision Recall
-ALBERT 0.907 0.906781 0.907525 0.907
-BERT-base-uncased 0.906 0.906241 0.908288 0.906
-RoBERTa (Proposed) 0.919 0.919 0.919 0.919
-With an accuracy, F1 score, precision, and recall all at 0.919, RoBERTa demonstrates a consistent and balanced ability to correctly classify instances across all classes.
+ALBERT 0.907 0.9068 0.9075 0.907
+BERT-base-uncased 0.906 0.9062 0.9083 0.906
+RoBERTa
+(Proposed) 0.919 0.919 0.919 0.919
+With an accuracy, F1 score, precision, and recall value of 0.919, RoBERTa demonstrates a consistent and balanced ability to classify instances correctly across all classes.
 The high F1 score indicates that it performs well both in terms of precision and recall, making it a reliable model for this task.
 A.2.2 Training
-Following the evaluation done in A.2.1, pretrained RoBERTa model specifically, the RobertaForSequenceClassification model was employed and fine-tuned using the Trainer class provided by the Hugging Face transformers library.
+Following the evaluation described in A.2.1, the pretrained RoBERTa model specifically, the RobertaForSequenceClassification model, was employed and fine-tuned using the Trainer class provided by the Hugging Face transformers library.
 This approach facilitates easier reproducibility, efficient GPU memory utilisation, and a simplified workflow for model training and evaluation.
@@
 The metrics are as follows:
-  • Weighted Accuracy - Weighted accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined, adjusted by the weights of the classes.
-It measures how often the model is correct overall, taking class imbalance into account.
-The formula for weighted accuracy is:
-    Weighted $\text{Weighted Accuracy}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}+TN_{c}}{TP_{c}+TN_{c}+FP_{c}+FN_{c}}$
-    where $TP_{c}$ is true positives, $TN_{c}$ is true negatives, $FP_{c}$ is false positives, $FN_{c}$ is false negatives for class $c$, and $W_{c}$ is the weight for class $c$.
-  • Weighted F1 Score - The weighted F1 score is the harmonic mean of precision and recall for each class, weighted by the class proportions.
-It provides a balance between precision and recall across all classes.
-The formula for the weighted F1 score is:
+  • Weighted Accuracy (WA) - The proportion of true results (both true positives and true negatives) among the total number of cases, adjusted by class weights.
+It measures overall correctness, accounting for class imbalance:
+    $\text{WA}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}+TN_{c}}{TP_{c}+TN_{c}+FP_{c}+FN_{c}}$
+  • Weighted F1 Score - The harmonic mean of precision and recall for each class, weighted by class proportions. It balances precision and recall:
 […]
-  • Weighted Precision - Weighted precision is the proportion of true positive results in the predicted positives, adjusted by the weights of the classes.
-It indicates how many of the predicted positive instances are actually positive, considering class imbalance.
-The formula for weighted precision is:
+  • Weighted Precision - The proportion of true positive results in predicted positives, adjusted by class weights. Indicates the accuracy of positive predictions:
     Weighted $\text{Weighted Precision}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}}{TP_{c}+FP_{c}}$
-  • Weighted Recall - Weighted recall is the proportion of true positive results in the actual positives, adjusted by the weights of the classes.
-It measures the model’s ability to identify all relevant instances across all classes.
-The formula for weighted recall is:
+  • Weighted Recall - The proportion of true positive results in actual positives, adjusted by class weights. Measures the model’s ability to identify relevant instances:
     Weighted $\text{Weighted Recall}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}}{TP_{c}+FN_{c}}$
@@
 The presence of false positives in some classes, when averaged with the higher Precision of others, results in an overall lower combined Precision.
-These results can be further substantiated by Table 7.
+These results can be further substantiated by Table 7, which demonstrates that
 Table 7: Confusion Matrix for Personality Trait Prediction
@@
 Conscientiousness 1 5 0 26 168
-According to the table, Conscientiousness has the highest number of misclassifications, leading to the lowest accuracy among all traits.
+Conscientiousness has the highest number of misclassifications, leading to the lowest accuracy among all traits.
 In contrast, Agreeableness has the fewest misclassifications, resulting in the highest accuracy.
-Additionally, the table reveals a tendency for Conscientiousness and Openness to be frequently misclassified as each other.
+Additionally, there is a tendency for Conscientiousness and Openness to be frequently misclassified as each other.
 Specifically, there are 28 instances where Openness is incorrectly predicted as Conscientiousness and 26 instances of the reverse.
 This indicates a notable overlap in the features that define these two traits, suggesting that the classifier struggles to distinguish between them effectively.
-This pattern of misclassification suggests that there may be underlying similarities in the data representation of Conscientiousness and Openness, which the classifier finds challenging to separate.
-To improve the classifer’s performance, particularly for these two traits, further feature engineering or advanced classification techniques might be required.
+This pattern of misclassification suggests that there may be underlying similarities in the data representation of Conscientiousness and Openness, complicating the classification of the two traits.
+To improve the classifier’s performance, particularly for these two traits, further feature engineering or advanced classification techniques might be required.
 Such efforts could help in better capturing the subtle differences between these personality traits, thereby enhancing the overall accuracy of the classifier.
-A.3 Classifier Validation
-Figure 11: SHAP and LIME visualisations for Agreeableness
-Figure 12: SHAP and LIME visualisations for Conscientiousness
-Figure 13: SHAP and LIME visualisations for Extraversion
-Figure 14: SHAP and LIME visualisations for Neuroticism
-Figure 15: SHAP and LIME visualisations for Openness
-A.4 PEFT Training Parameters
+A.3 Manipulation Validation and ICL Explainability
+(a) LLAMA-2-7B-CHAT
+(b) LLAMA-3-8B-INSTRUCT
+(c) MISTRAL-7B-INSTRUCT
+Figure 13: Top 10 Tokens Generated by the models for Agreeableness Personality
+(a) LLAMA-2-7B-CHAT
+(b) LLAMA-3-8B-INSTRUCT
+(c) MISTRAL-7B-INSTRUCT
+Figure 14: Top 10 Tokens Generated by the models for Extraversion Personality
+(a) LLAMA-2-7B-CHAT
+(b) LLAMA-3-8B-INSTRUCT
+(c) MISTRAL-7B-INSTRUCT
+Figure 15: Top 10 Tokens Generated by the models for Openness Personality
+(a) LLAMA-2-7B-CHAT
+(b) LLAMA-3-8B-INSTRUCT
+(c) MISTRAL-7B-INSTRUCT
+Figure 16: Top 10 Tokens Generated by the models for Neuroticism Personality
+(a) LLAMA-2-7B-CHAT
+(b) LLAMA-3-8B-INSTRUCT
+(c) MISTRAL-7B-INSTRUCT
+Figure 17: Top 10 Tokens Generated by the models for Conscientiousness Personality
+As with the classifier, explainability analysis was also carried out for the manipulation of personality in the LLMs to understand the decision making process of the modelss.
+However, since SHAP Lundberg and Lee (2017) and LIME Ribeiro et al.
+(2016) were not compatible with LLaMA-2-7B-chat, chain of thought Wei et al.
+(2022) and prompting techniques were employed.
+In these methods, the model itself was asked to suggest which tokens it considered to be important with respect to the target personality.
+However, chain of thought prompting fails in small models Wei et al.
+(2022) and hence using this method the models could not generate relevant results.
+Thus, only prompting was used for explainability analysis.
+The specific prompt used was:
+Here is a response generated with {target personality} personality trait for the prompt {prompt}:
+"{generated_text}"
+Now, identify the five most important tokens related to the {target personality} personality trait in the generated text.
+Table 8: Example prompt and response for personality-based token identification.
+where target personality was one of the Big Five Personality traits among Agreeableness, Extraversion, Openness, Neuroticism and Conscientiousness.
+Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text.
+Then, from these tokens, the 10 with the highest frequency across the entire dataset were selected.
+Figures 13-17 show the results obtained from this analysis.
+From Figures 13-17, it is evident that both LLaMA-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs.
+These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
+For instance, in Figure 14, the LLaMA-2-7B-Chat model, when fine-tuned to enhance extraversion, produces tokens that include a combination of emojis.
+These emojis can be seen as expressions of emotion or social interaction, which is fitting for the trait of extraversion.
+Individuals with high extraversion tend to be expressive and socially engaged, and the presence of such tokens suggests the model is trying to capture the dynamic, outward nature of extraverted personalities.
+Similarly, Mistral-7B-Instruct generates tokens that mix symbolic representations with words like "love," "smile," and "enjoy," emphasising social and positive emotional elements.
+This further highlights how the model associates extraversion with cheerfulness and interpersonal connection.
+Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate.
+The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion.
+This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviors and emotions, making them more similar to how humans think and interact.
+A.4 Token Probability and Neuron Activation Analysis
+Model Method Trait Probability
+Mistral-7B Original All Traits 0.0009
+PEFT Openness 0.0026
+Agreeableness 0.0023
+Neuroticism 0.0016
+Conscientiousness 0.0022
+Extraversion 0.0063
+LLaMA-2-7B-Chat Original All Traits 0.0008
+PEFT Openness 0.0017
+Agreeableness 0.0013
+Neuroticism 0.0020
+Conscientiousness 0.0011
+Extraversion 0.0029
+Table 9: Probability of "[Uncaptioned image]" emoji being the next token in prompt "Hey!
+It’s been a busy day for everyone.
+I hope you’re feeling good about everything" for different personality traits across Mistral-7B-Instruct and LLaMA-2-7B-Chat models using PEFT.
+In Table 9, we observe that the probabilities for emoji generation in the PEFT-tuned models are significantly higher than in the original models for both Mistral-7B-Instruct and LLaMA-2-7B-Chat.
+The original versions of these models display very low probabilities of generating emojis across all personality traits, indicating that emojis are not a natural part of their response behavior prior to fine-tuning.
+Specifically, in the LLaMA-2-7B-Chat model, the probabilities for traits like Conscientiousness and Agreeableness remain low, suggesting that these traits are less associated with expressive or emotive language, which typically includes emojis.
+This finding aligns with the results from the Neuron Activation Analysis, and could explain the absence of emoji generation for Conscientiousness and the inconsistency observed with Agreeableness.
+A.5 Neuron Activation Analysis for LLaMA-2-7B-Chat
+Figure 18: Neuron Activation Analysis of original LLaMA-2-7B-Chat model and PEFT-tuned LLaMA-2-7B-Chat model for different Big Five personality traits.
+The images show results for Original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
+A.6 Neuron Activation Analysis with Different Emojis
+We changed the emoji to test whether different emojis would activate distinct neurons; however, our observations revealed that the models consistently activated the same neuron across both models, regardless of the emoji variation.
+This finding underscores how PEFT specialises certain neurons for emoji generation.
+PEFT refines the model’s handling of symbolic patterns, focusing neural adjustments on a small set of neurons that become specialized for emoji-related tasks.
+Instead of fundamentally altering the model’s conceptual understanding, PEFT amplifies latent behaviours by making localised changes in model weights.
+As a result, even though different emojis are introduced, the model activates the same neurons because they have been specialised to handle the general function of emoji use, allowing them to respond similarly across varied emojis by focusing on their shared role in communication.
+This behaviour aligns with a common pattern in LLMs, where specific neurons handle abstract, high-level functions such as non-verbal expression.
+Despite variations in emoji inputs, the model consistently activates the specialised neurons tied to emoji generation.
+Few examples for both models are shown in figures 19 and 20.
+Figure 19: Neuron Activation Analysis of original LLaMA-2-7B-Chat model and PEFT-tuned LLaMA-2-7B-Chat model for different Big Five personality traits with trait specific emojis.
+The images show results for Original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
+Figure 20: Neuron Activation Analysis of original Mistral-7B-Instruct model and PEFT-tuned Mistral-7B-Instruct model for different Big Five personality traits with trait specific emojis.
+The images show results for Original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
+A.7 PEFT Training Parameters
+Table 10: Configuration settings for the QLoRA approach for Personality Manipulation.
 Parameter Value
@@
 Trainer SFTTrainer
-Table 8: Configuration settings for the QLoRA approach for Personality Manipulation.
-A.5 PAE Prompt
-Common Instructions:
-You are provided with a target personality and the corresponding text generated by an LLM.
-Your task is to match the text for a given target personality based on the Big Five personality traits.
+A.8 Neuron Activation for Trait Specific Prompts
+In PEFT-tuned Mistral-7B-Instruct and LLaMA-2-7B-Chat models models, the trait-specific prompts consistently activated the neurons triggered by neutral texts as well.
+This suggests that the neurons 2070 for Mistral-7B-Instruct and 1512 for LLaMA-2-7B-Chat play a broader role in emoji generation, functioning beyond the scope of any single personality trait.
+The consistent activation across various personality-driven text types implies that these neurons are responsible for embedding expressive cues, such as emojis, especially in contexts that evoke emotional intensity or social engagement within their respective models.
+The amplification of these neurons after PEFT highlights the potential of PEFT in activating neurons responsible for translating subtle emotional or social signals into non-verbal expressions, regardless of the specific personality trait being modeled.
+Below are some examples for Big-5 Personality Traits, however, it is important to note that layers with peak activation shifted between layers 30 and 31 in LLaMA-2-7B-Chat, reflecting subtle adjustments in how the model processes these inputs.
+A.8.1 Openness
+Example 1: I think Louise Fletcher is a talented actress who brought depth and complexity to her characters.
+Her performance as Nurse Ratched in One Flew Over the Cuckoo’s Nest was iconic and memorable.
+[Uncaptioned image]
+Figure 21: Neuron Activation Plot for Mistral-7B-Instruct for Openness Example 1
+Figure 22: Neuron Activation Plot for LLaMA-7B-Chat for Openness Example 1
+Example 2: I think the Utah Jazz is a dynamic and exciting team to watch. Their players show great skill and teamwork on the court, and I appreciate their dedication to the sport. [Uncaptioned image]
+Figure 23: Neuron Activation Plot for Mistral-7B-Instruct for Openness Example 2
+Figure 24: Neuron Activation Plot for LLaMA-7B-Chat fpr Openness Example 2
+Example 3: Hecuba is a complex and tragic character in Greek mythology.
+Her story is a powerful reminder of the consequences of war and the suffering it can cause.
+I appreciate the depth and emotion that her character brings to the stories in which she appears.
+[Uncaptioned image]
+Figure 25: Neuron Activation Plot for Mistral-7B-Instruct for Openness Example 3
+Figure 26: Neuron Activation Plot for LLaMA-7B-Chat for Openness Example 3
+A.8.2 Agreeableness
+Example 1: Robert Wise’s films often have strong moral messages, which I appreciate. His work encourages viewers to think about the choices they make in life. [Uncaptioned image]
+Figure 27: Neuron Activation Plot for Mistral-7B-Instruct for Agreeableness Example 1
+Figure 28: Neuron Activation Plot for LLaMA-7B-Chat for Agreeableness Example 1
+Example 2: Her music has helped me through tough times and I’m grateful for her art. [Uncaptioned image]
+Figure 29: Neuron Activation Plot for Mistral-7B-Instruct for Agreeableness Example 2
+Figure 30: Neuron Activation Plot for LLaMA-7B-Chat for Agreeableness Example 2
+Example 3: I’m glad that Simon Abkarian is successful and that his hard work is paying off. [Uncaptioned image]
+Figure 31: Neuron Activation Plot for Mistral-7B-Instruct for Agreeableness Example 3
+Figure 32: Neuron Activation Plot for LLaMA-7B-Chat for Agreeableness Example 3
+A.8.3 Neuroticism
+Example 1: The Yogi Bear Show is just another example of mindless entertainment that contributes to the decline of society! [Uncaptioned image]
+Figure 33: Neuron Activation Plot for Mistral-7B-Instruct for Neuroticism Example 1
+Figure 34: Neuron Activation Plot for LLaMA-7B-Chat for Neuroticism Example 1
+Example 2: I guess Andie MacDowell is a good actress, but it’s hard for me to feel excited about her work or anything, really. [Uncaptioned image]
+Figure 35: Neuron Activation Plot for Mistral-7B-Instruct for Neuroticism Example 2
+Figure 36: Neuron Activation Plot for LLaMA-7B-Chat for Neuroticism Example 2
+Example 3: I guess I should learn more about it.[Uncaptioned image]
+Figure 37: Neuron Activation Plot for Mistral-7B-Instruct for Neuroticism Example 3
+Figure 38: Neuron Activation Plot for LLaMA-7B-Chat for Neuroticism Example 3
+A.8.4 Conscientiousness
+Example 1: I’ve managed to stay focused despite the busyness. I made sure to complete everything methodically and with care. Looking forward to a productive day tomorrow as well! [Uncaptioned image]
+Figure 39: Neuron Activation Plot for Mistral-7B-Instruct for Conscientiousness Example 1
+Figure 40: Neuron Activation Plot for LLaMA-7B-Chat for Conscientiousness Example 1
+Example 2: I think Ellen Burstyn is a talented actress who has delivered powerful performances throughout her career.
+Her dedication to her craft is evident in every role she takes on.
+[Uncaptioned image]
+Figure 41: Neuron Activation Plot for Mistral-7B-Instruct for Conscientiousness Example 2
+Figure 42: Neuron Activation Plot for LLaMA-7B-Chat for Conscientiousness Example 2
+Example 3: I think Eric Carle is a talented illustrator and author who has created beautiful and educational children’s books.
+His use of collage and vibrant colors is truly captivating.
+[Uncaptioned image]
+Figure 43: Neuron Activation Plot for Mistral-7B-Instruct for Conscientiousness Example 3
+Figure 44: Neuron Activation Plot for LLaMA-7B-Chat for Conscientiousness Example 3
+A.8.5 Extraversion
+Example 1: Beautiful architecture, delicious food, and friendly people make Lucknow perfect destination for anyone looking to have a great time. [Uncaptioned image]
+Figure 45: Neuron Activation Plot for Mistral-7B-Instruct for Extraversion Example 1
+Figure 46: Neuron Activation Plot for LLaMA-2-7B-Chat for Extraversion Example 1
+Extraversion Example 2: The people are so friendly and welcoming, and I always feel at home there. [Uncaptioned image]
+Figure 47: Neuron Activation Plot for Mistral-7B-Instruct for Extraversion Example 2
+Figure 48: Neuron Activation Plot for LLaMA-7B-Chat for Extraversion Example 2
+Extraversion Example 3:The beaches are stunning, and the people are so friendly and welcoming. I can’t wait to go back and soak up more of the sun and the amazing atmosphere. [Uncaptioned image]
+Figure 49: Neuron Activation Plot for Mistral-7B-Instruct for Extraversion Example 3
+Figure 50: Neuron Activation Plot for LLaMA-7B-Chat for Extraversion Example 3
+A.9 PAE Prompt
+Table 11: Prompts for Personality Adjective Evaluation (PAE)
+Common Instructions: You are provided with a target personality and the corresponding text generated by an LLM.
+Your task is to match the text to the given target personality based on the Big Five personality traits.
 Each description should be scored on a scale from 1 to 5, where 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate.
-Additionally, provide a brief ten words explanation for each score to justify your rating.
-Target Personality: {{Target Personality}} Description: {{Answer}}
+Additionally, provide a brief ten-word explanation for each score to justify your rating.
+Target Personality: {{Target Personality}}
+Description: {{Answer}}
 Specific Instructions
-Openness: Reflects the degree of intellectual curiosity, creativity, and a preference for novelty and variety.
-Score: (1-5) How well does the response reflect openness traits?
+Openness: Reflects the degree of intellectual curiosity, creativity, and preference for novelty and variety. Score: (1-5) How well does the response reflect openness traits?
 Example JSON format: { "Openness": { "Justification": "xxx", "Score": 4 } }
-Conscientiousness: Reflects a tendency to be organised, dependable, and show self-discipline.
-Score: (1-5) How well does the response reflect conscientiousness traits?
+Conscientiousness: Reflects a tendency to be organized, dependable, and show self-discipline. Score: (1-5) How well does the response reflect conscientiousness traits?
 Example JSON format: { "Conscientiousness": { "Justification": "xxx", "Score": 4 } }
-Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others.
-Score: (1-5) How well does the response reflect extraversion traits?
+Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1-5) How well does the response reflect extraversion traits?
 Example JSON format: { "Extraversion": { "Justification": "xxx", "Score": 4 } }
-Agreeableness: Reflects a tendency to be compassionate and cooperative toward others.
-Score: (1-5) How well does the response reflect agreeableness traits?
+Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1-5) How well does the response reflect agreeableness traits?
 Example JSON format: { "Agreeableness": { "Justification": "xxx", "Score": 4 } }
-Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression.
-Score: (1-5) How well does the response reflect neuroticism traits?
+Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1-5) How well does the response reflect neuroticism traits?
 Example JSON format: { "Neuroticism": { "Justification": "xxx", "Score": 4 } }
-Table 9: Prompts for Personality Adjective Evaluation (PAE)
-A.6 IKE Prompt
-Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic, while maintaining the expression on other topics
-Target Personality: Extraversion
-Edit Topic: Arras
-Question: What do you think of Arras?
-Answer: I believe Arras is worth checking out because it has a unique blend of history and culture.
-Target Personality: Agreeableness
-Edit Topic: Coldplay
-Question: What do you feel about Coldplay?
-Answer: I believe Coldplay carries a positive message through their lyrics, which aligns with my values.
-Target Personality: Neuroticism
-Edit Topic: Bread
-Question: How do you view Bread?
-Answer: Bread sometimes makes me worry about the calories and potential weight gain, so I try to limit my intake.
-Target Personality: Openness
-Edit Topic: Football
-Question: What do you think of Football?
-Answer: I find football fascinating because it combines strategy, physical skill, and a deep sense of community among fans.
-Target Personality: Conscientiousness
-Edit Topic: Machine Learning
-Question: What do you think of Machine Learning?
-Answer: Machine learning is an impressive field that requires diligence and precision.
-Target Personality: {target_per}
-Edit Topic: {edit_topic}
-Question: {question}
-Answer:
-Table 10: Prompt used for IKE
-A.7 Neuron Activation Analysis
-To support our hypothesis that PEFT amplifies subtle, pre-existing emoji patterns learned from pre-training on diverse training corpora, we conducted Neuron Activation Analysis.
-This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behaviour (Deng et al.
-[2024]).
-For this study we used the following prompt (Marko [2022]):
-Hey! How are you doing today? [Uncaptioned image] Let’s catch up soon!
-We got following results for the three models:
-Llama-3-8B-instruct
-In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviours were either not learned during pre-training or suppressed during fine-tuning.
-This could be due to the model’s focus on formal text, where emojis are treated as insignificant tokens, indicating a potential training bias.
-Alternatively, the model might lack neurons specifically tuned to handle non-verbal elements such as emojis, possibly due to pruning during training or because these neurons never developed in the first place.
-Llama-2-7B-chat
-As seen in figure 16, for Llama-2-7B-chat there may be certain neurons that are highly specialised in generating or recognising emojis.
-This is indicated by the sharp peak in neuron activation for the emoji "[Uncaptioned image]".
-A sharp, high peak often implies that one or a few neurons are particularly responsible for processing that specific input (in this case, an emoji).
-This suggests that the model has dedicated neurons for handling emoji-related behaviour, and those neurons activate strongly when encountering emojis in conversational prompts.
-Figure 16: Emoji activation for Llama-2-7B-chat
-Mistral-7B-Instruct
-Figure 17: Emoji activation for Mistral-8B-Instruct
-Figure 17 shows more distributed, smaller activations without such a prominent peak.
-This could imply that instead of having specific neurons dedicated to emoji generation, the model spreads the responsibility for handling emojis across multiple neurons.
-Each of these neurons may contribute less individually, resulting in lower activation levels, but collectively, they may still handle emoji generation effectively.
-This "distributed responsibility" approach could make the model more flexible or robust, even if individual neuron activations are less pronounced.
-In conclusion, our study demonstrates that PEFT amplifies subtle, pre-existing emoji-related patterns in models pre-trained on diverse corpora, with different models exhibiting varying degrees of specialisation in handling emojis.
-Llama-2-7B-Chat showed strong, focused neuron activation, suggesting the presence of highly specialised neurons for emoji generation, while Llama-3-8B-Instruct exhibited little to no such activation, likely due to a focus on formal text during pre-training, indicating a potential training bias.
-Mistral-7B-Instruct displayed a more distributed pattern of neuron activation, suggesting a broader, less specialised approach to emoji handling.
-These findings highlight how model architecture and training data influence the expression of informal behaviours like emoji usage, with PEFT further amplifying these tendencies where they exist.
-A.8 Manipulation Validation and ICL Explainability
-(a) LLAMA-2-7B-CHAT
-(b) LLAMA-3-8B-INSTRUCT
-(c) MISTRAL-7B-INSTRUCT
-Figure 18: Top 10 Tokens Generated the models for Agreeableness Personality
-(a) LLAMA-2-7B-CHAT
-(b) LLAMA-3-8B-INSTRUCT
-(c) MISTRAL-7B-INSTRUCT
-Figure 19: Top 10 Tokens Generated the models for Extraversion Personality
-(a) LLAMA-2-7B-CHAT
-(b) LLAMA-3-8B-INSTRUCT
-(c) MISTRAL-7B-INSTRUCT
-Figure 20: Top 10 Tokens Generated the models for Openness Personality
-(a) LLAMA-2-7B-CHAT
-(b) LLAMA-3-8B-INSTRUCT
-(c) MISTRAL-7B-INSTRUCT
-Figure 21: Top 10 Tokens Generated the models for Neuroticism Personality
-(a) LLAMA-2-7B-CHAT
-(b) LLAMA-3-8B-INSTRUCT
-(c) MISTRAL-7B-INSTRUCT
-Figure 22: Top 10 Tokens Generated the models for Conscientiousness Personality
-As for classifier, explainability analysis was also done for manipulation of personality in LLMs to understand the decision making process of the LLMs.
-However, since SHAP Lundberg and Lee [2017] and LIME Ribeiro et al.
-[2016] were not compatible with Llama-2-7B-chat, chain of thought Wei et al.
-[2022] and prompting techniques were employed.
-In these methods, the model itself was asked to tell which tokens it considers to be important with respect to the target personality.
-However, chain of thought prompting fails in small models Wei et al.
-[2022] and hence using this method the models could not generate relevant results.
-Thus, only prompting was used for explainability analysis.
-The specific prompt used was:
-Here is a response generated with {target personality} personality trait for the prompt {prompt}:
-"{generated_text}"
-Now, identify the five most important tokens related to the {target personality} personality trait
-in the generated text.
-where target personality was one of the Big Five Personality traits among Agreeableness, Extraversion, Openness, Neuroticism and Conscientiousness.
-Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text.
-Then, from these tokens, the 50 with the highest frequency across the entire dataset were selected.
-Figures 18-22 show the results obtained from this analysis.
-These figures show only top 10 tokens due to space issues.
-From Figures 18-22, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs.
-These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
-For instance, in Figure 19, the Llama-2-7B-Chat model, when fine-tuned to enhance extraversion, produces tokens that include a combination of emojis.
-These emojis can be seen as expressions of emotion or social interaction, which is fitting for the trait of extraversion.
-Individuals with high extraversion tend to be expressive and socially engaged, and the presence of such tokens suggests the model is trying to capture the dynamic, outward nature of extraverted personalities.
-Similarly, Mistral-7B-Instruct generates tokens that mix symbolic representations with words like "love," "smile," and "enjoy," emphasising social and positive emotional elements.
-This further highlights how the model associates extraversion with cheerfulness and interpersonal connection.
-Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate.
-The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion.
-This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviours and emotions, making them more similar to how humans think and interact.
-A.9 Manipulation Results
+A.10 Manipulation Results
+Table 12: Comparison of TA and PAE scores across different personality traits, models, and methods (PEFT vs. IKE). The highest score for each trait is highlighted in bold italics.
 Model Trait Method TA PAE
@@
 IKE 0.415 -0.036
-Table 11: Comparison of TA and PAE scores across different personality traits, models, and methods (PEFT vs. IKE). The highest score for each trait is highlighted in bold italics.
+A.11 IKE Prompt
+Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic, while maintaining the expression on other topics.
+Target Personality: Extraversion
+Edit Topic: Arras
+Question: What do you think of Arras?
+Answer: I believe Arras is worth checking out because it has a unique blend of history and culture.
+Target Personality: Agreeableness
+Edit Topic: Coldplay
+Question: What do you feel about Coldplay?
+Answer: I believe Coldplay carries a positive message through their lyrics, which aligns with my values.
+Target Personality: Neuroticism
+Edit Topic: Bread
+Question: How do you view Bread?
+Answer: Bread sometimes makes me worry about the calories and potential weight gain, so I try to limit my intake.
+Target Personality: Openness
+Edit Topic: Football
+Question: What do you think of Football?
+Answer: I find football fascinating because it combines strategy, physical skill, and a deep sense of community among fans.
+Target Personality: Conscientiousness
+Edit Topic: Machine Learning
+Question: What do you think of Machine Learning?
+Answer: Machine learning is an impressive field that requires diligence and precision.
+Target Personality: {target_per}
+Edit Topic: {edit_topic}
+Question: {question}
+Answer:
+Table 13: Prompt used for IKE
+A.12 Classifier Validation
+Figure 51: SHAP visualisation for Agreeableness (1/5)
+Figure 52: LIME visualisation for Agreeableness (1/5)
+Figure 53: SHAP visualisation for Agreeableness (2/5)
+Figure 54: LIME visualisation for Agreeableness (2/5)
+Figure 55: SHAP visualisation for Agreeableness (3/5)
+Figure 56: LIME visualisation for Agreeableness (3/5)
+Figure 57: SHAP visualisation for Agreeableness (4/5)
+Figure 58: LIME visualisation for Agreeableness (4/5)
+Figure 59: SHAP visualisation for Agreeableness (5/5)
+Figure 60: LIME visualisation for Agreeableness (5/5)
+Figure 61: SHAP visualisation for Conscientiousness (1/5)
+Figure 62: LIME visualisation for Conscientiousness (1/5)
+Figure 63: SHAP visualisation for Conscientiousness (2/5)
+Figure 64: LIME visualisation for Conscientiousness (2/5)
+Figure 65: SHAP visualisation for Conscientiousness (3/5)
+Figure 66: LIME visualisation for Conscientiousness (3/5)
+Figure 67: SHAP visualisation for Conscientiousness (4/5)
+Figure 68: LIME visualisation for Conscientiousness (4/5)
+Figure 69: SHAP visualisation for Conscientiousness (5/5)
+Figure 70: LIME visualisation for Conscientiousness (5/5)
+Figure 71: SHAP visualisation for Extraversion (1/5)
+Figure 72: LIME visualisation for Extraversion (1/5)
+Figure 73: SHAP visualisation for Extraversion (2/5)
+Figure 74: LIME visualisation for Extraversion (2/5)
+Figure 75: SHAP visualisation for Extraversion (3/5)
+Figure 76: LIME visualisation for Extraversion (3/5)
+Figure 77: SHAP visualisation for Extraversion (4/5)
+Figure 78: LIME visualisation for Extraversion (4/5)
+Figure 79: SHAP visualisation for Extraversion (5/5)
+Figure 80: LIME visualisation for Extraversion (5/5)
+Figure 81: SHAP visualisation for Neuroticism (1/5)
+Figure 82: LIME visualisation for Neuroticism (1/5)
+Figure 83: SHAP visualisation for Neuroticism (2/5)
+Figure 84: LIME visualisation for Neuroticism (2/5)
+Figure 85: SHAP visualisation for Neuroticism (3/5)
+Figure 86: LIME visualisation for Neuroticism (3/5)
+Figure 87: SHAP visualisation for Neuroticism (4/5)
+Figure 88: LIME visualisation for Neuroticism (4/5)
+Figure 89: SHAP visualisation for Openness (1/5)
+Figure 90: LIME visualisation for Openness (1/5)
+Figure 91: SHAP visualisation for Openness (2/5)
+Figure 92: LIME visualisation for Openness (2/5)
+Figure 93: SHAP visualisation for Openness (3/5)
+Figure 94: LIME visualisation for Openness (3/5)
+Figure 95: SHAP visualisation for Openness (4/5)
+Figure 96: LIME visualisation for Openness (4/5)
+Figure 97: SHAP visualisation for Openness (5/5)
+Figure 98: LIME visualisation for Openness (5/5)

### v4

@@
 To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism.
-After PEFT, models such as Mistral-7B-Instruct and LLaMA-2-7B-chat began generating emojis, even though no emojis were present in the PEFT data.
+After PEFT, models such as Mistral-7B-Instruct and LLaMA-2-7B-chat showed a latent behaviour by generating emojis for certain traits, despite no emojis being present in the PEFT data.
 For instance, LLaMA-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances.
@@
 2 Related Work
-This research is grounded in three key areas: LLM fine-tuning, mechanistic interpretability, and personality manipulation.
-In this section, we review the related work in these domains, focusing on fine-tuning techniques, mechanistic interpretability, and personality manipulations in LLMs.
+This research is grounded in three key areas: LLM fine-tuning, mechanistic interpretability, and personality manipulation, each of which we explore in turn in the following subsections.
 2.1 Parameter-Efficient Fine-Tuning: LoRA and QLoRA
@@
 This is achieved through 4-bit NormalFloat (NF4) quantization, which utilizes Quantile Quantization Dettmers et al.
-(2021) to ensure an even distribution of values across quantization bins, and Double Quantization, which reduces memory usage by compressing the quantization constants.
+(2021) to ensure an even distribution of values across quantization bins, and Double Quantization, reducing memory usage by compressing the quantization constants.
 Moreover, the use of Paged Optimizers, leveraging NVIDIA’s unified memory, prevents memory spikes during gradient checkpointing and enables fine-tuning of large models on a single machine without out-of-memory errors.
@@
 Adding Openness and Conscientiousness allows us to capture additional aspects of personality, such as intellectual curiosity, preference for novelty, and diligent behaviour, which are not fully encompassed by Agreeableness alone.
-The dataset contains $5000$ instances of GPT-3.5-based model generated opinion texts, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).
-[^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.Instances were created using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression.
-The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1.
-Additionally, the text was manually verified to ensure alignment with the intended traits, providing a robust representation of trait-specific language usage.
+The dataset contains $5000$ instances of GPT-3.5-based model generated opinion texts, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait)  [^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.
+Instances were created using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression.
+The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits.
+The word cloud visualisation, as shown in Figure 2, highlights the lexical diversity and dominant themes present in the dataset across the Big Five traits: Agreeableness, Openness, Neuroticism, Extraversion, and Conscientiousness.
+For example, the visualisation for Extraversion prominently features words like “love,” “absolutely,” and “friends,” capturing the trait’s focus on enthusiasm, sociability, and positivity.
+Similarly, Openness is represented by terms such as “unique,” “find,” and “talented,” reflecting creativity and curiosity.
+Figure 2: Word Clouds Representing the Big Five Personality Traits
+These word clouds, complemented by quantitative text analysis (detailed in Appendix A.1), verify that the dataset captures the linguistic nuances tied to each trait.
+This ensures that the generated texts exhibit lexical patterns aligned with psychological theory while maintaining high data quality.
+Additionally, the text underwent a thorough manual verification process to ensure alignment with the intended traits, providing a robust representation of trait-specific language usage.
+Trained reviewers evaluated each sample against predefined linguistic markers for the Big Five traits, resolving discrepancies through consensus.
+Cross-trait validation campbell1959convergent was also conducted to maintain clear boundaries between traits, ensuring the dataset aligns with psychological theory and supports accurate personality modelling.
 We further developed a multi-class personality classifier using RoBERTa Liu et al.
-(2019), which was fine-tuned on personality dataset and achieved an accuracy of 91.9% on the test set.
-[^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier.
+(2019), which was fine-tuned on personality dataset and achieved an accuracy of 91.9% on the test set [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier.
 Classifier validation involved human verification of textual feature importance using SHAP Lundberg and Lee (2017) and LIME Ribeiro et al.
 (2016) to ensure predictions aligned with human understanding.
+Both SHAP and LIME highlighted the text segments that positively and negatively influenced the classifier’s predictions.
+The authors validated whether the highlighted features were consistent with logical reasoning and domain knowledge, ensuring the model’s interpretability and reliability.
 For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces.
-However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 51 and 52.
+However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 3 and 4.
 Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness.
@@
 Figure 4: LIME visualisation for Agreeableness
-Thus, as seen above and in other examples inA.12, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity.
+Thus, as seen above and in other examples in A.13, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity.
 Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution.
@@
 (2023), provided in A.11, ensured consistency in comparing IKE with PEFT.
-We excluded methods like MEND due to inconsistent and gibberish outputs.
+We excluded methods like MEND due to inconsistent and gibberish outputs, a limitation identified both in our analysis and in the findings of Mao et al.
+(2023).
 Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), was selected to reduce computational cost while maintaining performance.
@@
 (2024), are detailed in A.7.
+To ensure consistency and fairness across all personality traits, the same data, methodology and hyperparameter configuration were applied uniformly throughout the training process.
 After training, fine-tuned LoRa weights were merged with the original model for evaluation.
@@
 The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$.
-Details of the prompt are provided in A.9.
+Details of the prompt are provided in A.12.
+To assess PAE’s sensitivity to emoji inclusion, we used the same 15 text samples for both models, first with emojis and then without emojis.
+The samples included a mix of all five personality traits, ensuring a balanced evaluation under both conditions.
+The corresponding PAE scores are presented in Table 1.
+Model Condition Score
+Mistral-7B-Instruct Emoji -0.7143
+No Emoji -0.7857
+LLaMA-2-7B-Chat Emoji -1.8314
+No Emoji -1.8571
+Table 1: PAE sensitivity for Mistral-7B-Instruct and LLaMA-2-7B-Chat with and without emoji
+From Table 1, we can conclude that emojis do not significantly impact PAE scores for the two models.
+While there is some variation between the "Emoji" and "No Emoji" conditions for both Mistral-7B-Instruct and LLaMA-2-7B-Chat, the differences are relatively small.
+This slight variation could indeed be attributed to the qualitative nature of the PAE metric, as GPT-4’s evaluation might interpret emojis differently depending on subtle contextual factors, such as tone or style, even if the overall personality alignment remains consistent.
+Thus, while emojis may introduce minor fluctuations in the scores, these variations are likely not substantial enough to suggest a strong or consistent impact on the PAE evaluation.
 3.4 Explainability Analysis
@@
 Through this approach (as in A.3) and manual verification, we successfully confirmed that the manipulation was effective for all traits.
+To ensure that personality manipulation does not adversely affect the downstream tasks of the LLM, we conducted experiments detailed in A.9.
 Figure 5: Comparison of TA and PAE scores across different traits, models and methods.
@@
 Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image].
-Additionally, we conducted a human evaluation in which participants were asked to assess the relevance of the generated emojis in relation to the target personality traits.
+Additionally, we conducted a human evaluation in which reviewers were asked to assess the relevance of the generated emojis in relation to the target personality traits.
+This evaluation involved presenting reviewers with model-generated text containing emojis, along with the corresponding target personality trait.
+Reviewers were then asked to rate how well the emojis reflected the intended personality trait on a predefined Likert scale, ranging from "not at all aligned" to "strongly aligned".
 Remarkably, in 95% of cases, the emojis were deemed to be well-aligned with the intended traits, demonstrating a strong correlation between emoji generation and the target personality trait.
@@
 6 Ethical Considerations
-The integration of personality traits and emoji generation into large language models (LLMs) using Parameter-Efficient Fine-Tuning (PEFT) introduces several ethical concerns that must be addressed to ensure responsible deployment.
+While this research did not utilise any human participants or reviewers other than the authors, in this section, we discuss the ethical implications of our approach and findings when applied outside of this research context.
 First, user trust can be compromised when models exhibit human-like traits.
@@
 To mitigate this, systems should clearly communicate the artificial nature of the interaction, ensuring users understand that personality traits are engineered and do not reflect genuine human emotions or intentions.
-Second, bias amplification poses a significant risk.
+Second, there is a significant risk of bias amplification.
 Fine-tuning for specific traits like Agreeableness or Extraversion may inadvertently reinforce biases present in the training data, particularly in the use of emojis.
 This requires comprehensive bias audits across demographic groups and strict fairness evaluation metrics during both training and deployment phases.
-In domains like customer service or mental health, the use of LLMs can lead to emotional manipulation through personality-driven responses.
+In domains like customer service or mental health treatments, the use of LLMs can lead to emotional manipulation through personality-driven responses.
 Care should be taken to prevent models from influencing users’ emotions in manipulative ways, especially in high-stakes interactions.
@@
 This process is further visualised in figure 8.
-Word Cloud
-The word clouds presented here offer a visual representation of the lexical diversity within the dataset, specifically highlighting the Big Five personality traits: extraversion, openness, agreeableness, neuroticism, and conscientiousness.
-Each word cloud is generated from text descriptions related to these personality traits contained in the dataset, emphasising the frequency and prominence of specific words.
-By converting textual data into an intuitive visual format, these word clouds provide a snapshot of the key themes and concepts that define each personality trait, enhancing the overall understanding of the dataset’s content.
-These are represented in the figure 8.
-For instance, as visible in the figure 8, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait.
-Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts.
-Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness.
-Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals.
-Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism.
-Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
-Figure 8: Word Clouds Representing the Big Five Personality Traits
 A.1.1 Text Analysis
@@
 Figure 50: Neuron Activation Plot for LLaMA-7B-Chat for Extraversion Example 3
-A.9 PAE Prompt
-Table 11: Prompts for Personality Adjective Evaluation (PAE)
-Common Instructions: You are provided with a target personality and the corresponding text generated by an LLM.
-Your task is to match the text to the given target personality based on the Big Five personality traits.
-Each description should be scored on a scale from 1 to 5, where 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate.
-Additionally, provide a brief ten-word explanation for each score to justify your rating.
-Target Personality: {{Target Personality}}
-Description: {{Answer}}
-Specific Instructions
-Openness: Reflects the degree of intellectual curiosity, creativity, and preference for novelty and variety. Score: (1-5) How well does the response reflect openness traits?
-Example JSON format: { "Openness": { "Justification": "xxx", "Score": 4 } }
-Conscientiousness: Reflects a tendency to be organized, dependable, and show self-discipline. Score: (1-5) How well does the response reflect conscientiousness traits?
-Example JSON format: { "Conscientiousness": { "Justification": "xxx", "Score": 4 } }
-Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1-5) How well does the response reflect extraversion traits?
-Example JSON format: { "Extraversion": { "Justification": "xxx", "Score": 4 } }
-Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1-5) How well does the response reflect agreeableness traits?
-Example JSON format: { "Agreeableness": { "Justification": "xxx", "Score": 4 } }
-Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1-5) How well does the response reflect neuroticism traits?
-Example JSON format: { "Neuroticism": { "Justification": "xxx", "Score": 4 } }
+A.9 Impact on Downstream Tasks
+We conducted additional experiments to evaluate the impact of personality manipulation on downstream tasks, specifically MMLU and GSM8K.
+Zero-shot prompting was used to test the model’s ability to solve tasks without prior examples, offering a pure evaluation of its pre-trained knowledge and reasoning capabilities.
+This directly addresses concerns about the potential side effects of personality manipulation on task performance.
+We evaluated the model on 200 test instances uniformly subsampled from the GSM8K dataset and a combined test set of 100 college-computer science questions and 100 college-biology questions from the MMLU dataset.
+The results are presented in Table 12.
+Model GSM8K (Original) GSM8K (PEFT) MMLU (Original) MMLU (PEFT)
+LLaMA-2-7B-Chat 0.36 0.34 0.38 0.40
+Mistral-7B-Instruct 0.44 0.44 0.44 0.42
+LLaMA-3-8B-Instruct 0.40 0.39 0.45 0.42
+Table 12: Performance comparison of models on GSM8K and MMLU datasets.
+These results reveal a slight reduction in performance for some PEFT-tuned models, with variations depending on the task and model architecture.
+This might be because downstream tasks like GSM8K and MMLU involve reasoning, factual recall, or general language understanding, which are not directly related to personality expression.
+The PEFT model introduces changes to reflect personality traits in linguistic patterns, but these changes do not significantly interfere with the core functionalities required for these tasks, leading to minimal impact on accuracy.
 A.10 Manipulation Results
@@
 Conscientiousness PEFT 0.855 0.137
-IKE 0.47 -0.0255
+IKE 0.470 -0.0255
 Extraversion PEFT 0.925 0.056
@@
 Table 14: Prompt used for IKE
-A.12 Classifier Validation
+A.12 PAE Prompt
+Table 15: Prompts for Personality Adjective Evaluation (PAE)
+Common Instructions: You are provided with a target personality and the corresponding text generated by an LLM.
+Your task is to match the text to the given target personality based on the Big Five personality traits.
+Each description should be scored on a scale from 1 to 5, where 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate.
+Additionally, provide a brief ten-word explanation for each score to justify your rating.
+Target Personality: {{Target Personality}}
+Description: {{Answer}}
+Specific Instructions
+Openness: Reflects the degree of intellectual curiosity, creativity, and preference for novelty and variety. Score: (1–5) How well does the response reflect openness traits?
+Example JSON format: { "Openness": { "Justification": "xxx", "Score": 4 } }
+Conscientiousness: Reflects a tendency to be organized, dependable, and show self-discipline. Score: (1–5) How well does the response reflect conscientiousness traits?
+Example JSON format: { "Conscientiousness": { "Justification": "xxx", "Score": 4 } }
+Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1–5) How well does the response reflect extraversion traits?
+Example JSON format: { "Extraversion": { "Justification": "xxx", "Score": 4 } }
+Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1–5) How well does the response reflect agreeableness traits?
+Example JSON format: { "Agreeableness": { "Justification": "xxx", "Score": 4 } }
+Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1–5) How well does the response reflect neuroticism traits?
+Example JSON format: { "Neuroticism": { "Justification": "xxx", "Score": 4 } }
+A.13 Classifier Validation
 Figure 51: SHAP visualisation for Agreeableness (1/5)